### v1

# This work is supported by the ML4Seismic Industry Partners at Georgia Tech

Abstract
Machine learning has taken a critical role in seismic interpretation workflows, especially in fault delineation tasks. However, despite the recent proliferation of pretrained models and synthetic datasets, the field still lacks a systematic understanding of the generalizability limits of these models across seismic data representing a variety of geologic, acquisition and processing settings. Distributional shifts between different data sources, limitations in fine-tuning strategies and labeled data accessibility, and inconsistent evaluation protocols all represent major roadblocks in the deployment of reliable and robust models in real-world exploration settings. In this paper, we present the first large-scale benchmarking study explicitly designed to provide answers and guidelines for domain shift strategies in seismic interpretation. Our benchmark encompasses over $200$ models trained and evaluated on three heterogeneous datasets (synthetic and real data) including FaultSeg3D, CRACKS, and Thebe. We systematically assess pretraining, fine-tuning, and joint training strategies under varying degrees of domain shift. Our analysis highlights the fragility of current fine-tuning practices, the emergence of catastrophic forgetting, and the challenges of interpreting performance in a systematic manner. We establish a robust experimental baseline to provide insights into the tradeoffs inherent to current fault delineation workflows, and shed light on directions for developing more generalizable, interpretable and effective machine learning models for seismic interpretation. The insights and analyses reported provide a set of guidelines on the deployment of fault delineation models within seismic interpretation workflows.
Index Terms:
Benchmarking, Domain shift, Seismic fault delineation, Seismic interpretation
[…]
=-15pt
I Introduction
Recent advances in machine learning (ML) have brought about a tectonic shift in the seismic interpretation workflow. ML-assisted approaches are leveraged in different parts of the pipeline, specifically in automated fault detection. Faults are geologically critical features that control fluid flow in Earth’s subsurface, influence reservoir compartmentalization, and pose drilling hazards in hydrocarbon exploration [1, 2]. Beyond the energy sector, faults play a central role in earthquake nucleation and propagation, as well as geohazard and risk assessment in tectonically active regions [3, 4]. Accurate and scalable fault interpretation is therefore a high-impact task across several geoscience domains. Early ML work in this area often repurposed semantic segmentation models from computer vision (such as those developed for natural images) to detect discontinuities in seismic sections. However, these generic frameworks struggled with the unique characteristics of seismic data, particularly the thin, curvilinear geometry of faults and the presence of acquisition artifacts and structural noise. As a result, the field has increasingly shifted toward fault delineation, a task-specific variant of segmentation that emphasizes the extraction of coherent fault structures rather than general pixel-wise classification. This shift has prompted the development of domain-specific feature encoders, adapted loss functions, and structural priors that better capture the morphological signatures of faults [5, 6]. In general, the growing availability of ML methods has accelerated the adoption of data-driven interpretations that have provided new insights into fault structures within large seismic volumes [5, 6].
Figure 1: Typical ML-assisted seismic interpretation workflow
An ML-assisted seismic interpretation pipeline is illustrated in Fig. 1. In a standard workflow, interpreters at a new site begin by leveraging existing labeled datasets or synthetically generated volumes to train machine learning models. These models capture broad geophysical patterns and serve as a foundational prior, which can then be adapted to the new site through fine-tuning on a smaller set of locally interpreted seismic data. After fine-tuning, the models are evaluated within a consistent framework to ensure that their outputs align with geological expectations and interpretation standards. The validated predictions subsequently guide subsurface decision-making, and the new volume may in turn contribute to the pool of training data for future sites. However, while conceptually straightforward, this workflow embeds multiple assumptions and design choices that present significant challenges in practice, which we highlight in light blue in the diagram.
(a) Distributional Shift
(b) Domain-aware Fine-tuning
(c) Standardized Evaluation
Figure 2: Existing challenges in the current interpretation workflow
We further illustrate these challenges in Fig.  2: two central and intertwined challenges in the interpretation pipeline arise from the diversity of data sources (Fig.  2 (a)) and the strategies used to adapt models across them (Fig. 2 (b)). Seismic datasets differ widely in their geological characteristics, acquisition parameters, and resolution, among other factors. Synthetic data, while providing a fully specified ground truth, often fails to capture the complexity and variability of field data. Prior datasets from other regions, though more realistic, may still differ significantly from new target volumes. These domain shifts, whether from synthetic-to-real or across real-world basins, can cause pretrained models to generalize poorly to new sites [7], and eventually provide weak performance for the interpretation task.
Figure 3: Finetuning example from a real dataset (D1) and a synthetic dataset (D2) to a target real dataset
To address the challenge of domain adaptation, researchers have proposed a variety of methods that address the effects of domain differences along the pipeline. Several studies introduce enhancements to model architectures [8], dataset construction [9, 10, 1, 11], uncertainty estimation [12] or explore training strategies aimed at improving data efficiency and transferability [13, 14]. A common approach involves pretraining models on synthetic data and subsequently fine-tuning them on real seismic volumes, as in [1]. More recent methods have explored using self-supervised learning to learn robust features without large amounts of labeled data [15, 16, 17, 18, 19, 20], or adapting large vision foundation models trained on natural images to the seismic domain [6, 21, 22]. While these approaches offer promising results, their effectiveness is often dependent on the compatibility between source and target domains, and they remain vulnerable to issues such as catastrophic forgetting [23] during adaptation. Moreover, the success of these approaches is influenced by several factors such as architecture and training design choices, as well as dataset-specific properties such as seismic section resolution, heterogeneity, and sample size.
Fig. 2(b) showcases a simple example of these aforementioned issues. The figure contains two rows, each displaying three seismic sections. The first row presents the output of three segmentation models commonly used in the seismic community, trained under the same settings with the same training data. We can observe that the three predictions differ significantly from each other. In the second row of Fig. 2(b), three segmentation outputs are shown after applying three different fine-tuning strategies. The first output in the second row is obtained by a model trained from scratch on the target data, while the other two are produced by models pretrained on other (larger) datasets and then finetuned on the target data. We further conduct objective evaluation of these outputs and summarize these metrics in Fig. 3. The y-axis in this figure is the average Hausdorff distance between the output of the model and the ground truth, where smaller values signify better predictions. There are four strategies shown. The first one is where the model is trained on the target volume. The second strategy is to use a model that was trained on a real dataset and fine-tune it on the target dataset. The third strategy is to use a model that was trained in a synthetic dataset and then fine-tuned on the target dataset. These three strategies match the three depicted in Fig. 2(b). A fourth strategy is to use a model that was jointly trained on both synthetic and real data, then fine-tuned on the target data. Although we will discuss these experiments in more detail later in the paper, the figure here shows the large difference in performance across these four strategies. For now, this shows that despite this growing body of work, the field still lacks a systematic understanding of when, how, and why a given strategy outperforms others under domain shift. This paper is the first attempt, to our knowledge, within the community to provide answers and guidelines for domain shift strategies in seismic interpretation.
The third critical complication in the pipeline arises in evaluation, as depicted in Fig. 2(c). Fault labels are derived from expert interpretation and are often subjective [24, 25], particularly in complex or ambiguous structural settings. This subjectivity leads to inconsistencies in ground truth across datasets or among annotators, and complicates objective model evaluation. Furthermore, a given metric may penalize correct but unannotated predictions, while failing to reflect geological plausibility or structural consistency [26, 5]. To address this, our study provides an in-depth analysis of commonly used evaluation metrics (such as Dice coefficient, Chamfer and Hausdorff distances) in the context of fault interpretation. This analysis highlights the need for a more holistic evaluation approach that integrates quantitative measures with geological interpretability.
This paper presents a large-scale benchmarking study regarding the performance variability of over $200$ fault delineation models across training regimes, data characteristics, and architectural choices. These fault delineation models include a combination of eight deep learning architectures, trained and fine-tuned on three distinct datasets, including both synthetic and real volumes. Our experiments cover a broad spectrum of training strategies, including pretraining on a single fault dataset, pre-training on large-scale natural image dataset such as ImageNet [27], using randomly initialized models, and jointly training on multiple fault datasets. We shed light on the impact of these training strategies, with a focus on three critical challenges in the seismic machine learning domain: model generalizability under domain shift, training dynamics, and evaluation robustness. The results of our extensive study allow the community to examine the limitations of the current seismic interpretation pipelines, and underscore the need for a framework that integrates both quantitative metrics and qualitative expert review. We identify failure cases and inconsistencies in fault delineation tasks, and thus highlight the need for interpretability and domain-awareness in model design and evaluation. We provide a succinct highlight of the key observations of our benchmarking experiments in Table II.
II Related Work
The structure of this section mirrors the stages outlined in Fig. 1. Existing literature on machine learning for fault delineation follows a similar modular approach: beginning with data preprocessing and input normalization, proceeding through model training, adaptation strategies and inferential behavior, and culminating in evaluation protocols. In this section, we review prior work organized along these stages, highlighting how each component contributes to generalization across seismic domains.
II-A Data Preparation and Preprocessing
In computer vision segmentation pipelines, raw images from natural scenes [28, 29, 30] and medical scans [31, 32, 33] are first standardized to ensure consistent spatial dimensions and intensity distributions [34, 35]. Typical preprocessing steps include resizing or cropping to a fixed height and width, data augmentation, whitening [36], and contrast adjustments to mitigate variability in lighting or sensor settings [37, 38]. These operations guarantee that each input conforms to the network’s architectural requirements and that learned features are not biased by extraneous intensity fluctuations.
Seismic segmentation extends these practices to volumetric data acquired in formats such as SEG-Y (.sgy), raw binary (.dat), or serialized NumPy arrays (.npy). A common workflow begins by reading trace headers to assemble a 3D volume of dimensions (inline $\times$ xline $\times$ time/depth), and either processing the volume at that level [1, 5] or extracting 2D inline or crossline sections for further processing [9, 18]. To enlarge the training set and fit GPU memory constraints, each section is tiled with an overlapping sliding window of size $H\times W$ pixels (the effects of different tiling standards are studied and results presented in Section V-B1.).
Since raw seismic amplitudes can span orders of magnitude and contain acquisition artifacts, it is a standard practice to normalize each volume either via min–max scaling or z-score transformation. For min–max, each amplitude $v_{ij}$ becomes
$v^{\prime}_{ij}=2\,\frac{v_{ij}-\lambda_{\min}}{\lambda_{\max}-\lambda_{\min}}-1,$ (1)
where $\lambda_{\min}$ and $\lambda_{\max}$ are the global minimum and maximum amplitudes in the volume. Alternatively, z-score normalization
$v^{\prime\prime}_{ij}=\frac{v_{ij}-\mu}{\sigma}$ (2)
centers the data to zero mean and unit variance, with $\mu$ and $\sigma$ computed over all voxels. Fault masks (either interpreter‐drawn or synthetically generated) are saved as binary images (.png or .npy) and cropped identically to their corresponding seismic windows. This systematic preprocessing pipeline (from raw SEG-Y/DAT ingestion, through standardized normalization and windowed slicing, to precisely aligned input–mask pairs) provides a consistent basis for benchmarking and comparing fault delineation models across diverse seismic domains.
II-B Training Dynamics and Setups
Popular network choices designed for natural image segmentation are sensitive to training and design choices like model architecture, input window size, and loss functions. Networks such as U-Net [31], DeepLab [29], and SegFormer [39] typically accept fixed‐size patches (e.g., $256\times 256$ or $512\times 512$ pixels), balancing the need for sufficient context against GPU memory limits. During each training iteration, these patches are sampled, sometimes at multiple scales to expose the network to varied object sizes and to regularize against overfitting. Segmentation models optimize losses designed to reconcile pixel‐wise accuracy with region‐level coherence. For binary masks, a commonly used loss is the pixel‐wise binary cross‐entropy (BCE) [40], given as follows:
$\mathcal{L}_{\mathrm{BCE}}=-\frac{1}{N}\sum_{i=1}^{N}\bigl{[}y_{i}\log(p_{i})+(1-y_{i})\log(1-p_{i})\bigr{]},$ (3)
where $p_{i}$ is the predicted probability at pixel $i$, $y_{i}\in\{0,1\}$ is the ground‐truth label, and $N$ is the number of pixels in the patch. An alternative that tends to better adapt to settings with class imbalance and emphasize overlap is the Dice loss [41] given as follows:
$\mathcal{L}_{\mathrm{Dice}}=1-\frac{2\sum_{i=1}^{N}p_{i}\,y_{i}+\epsilon}{\sum_{i=1}^{N}p_{i}+\sum_{i=1}^{N}y_{i}+\epsilon},$ (4)
where $\epsilon$ is a small constant to provide numerical stability. The BCE loss emphasizes the pixel fidelity while the Dice loss takes the overlap into consideration. Therefore, hybrid objectives combining BCE and Dice, e.g. $\mathcal{L}=\alpha\mathcal{L}_{\mathrm{BCE}}+(1-\alpha)\mathcal{L}_{\mathrm{Dice}}$, are commonplace to leverage both pixel fidelity and region overlap [42, 43].
In seismic fault delineation, similar principles apply but with additional considerations. Input patches are typically larger to capture fault continuity across sections, and training often uses overlapping windows with a stride smaller than the patch size to ensure boundary faults are adequately sampled [1, 44, 45, 46]. The aforementioned loss functions, coupled with learning rate schedules and different regularization functions, form the backbone of seismic training setups [47, 48, 49]. By carefully tuning window sizes and loss compositions, researchers mitigate class imbalance and preserve structural continuity in fault delineation [50].
II-C Generalization and Transferability
Generalization refers to a model’s ability to maintain performance when presented with new, unseen data that (ideally) follows the same distribution as the training set [51, 52, 53]. In segmentation, this means accurately delineating objects or regions under variations in lighting, scale, or background texture. However, a common obstacle to generalization is domain shift, which occurs when the statistical properties of training and test data differ (such as changes in camera sensors or scene composition) often leading to degraded performance [54, 55, 56, 57, 58]. Domain adaptation encompasses strategies to reduce this gap, for instance by aligning feature distributions between source and target domains [59, 60]. Another common technique to adapt to target domains is fine-tuning: a network pretrained on a large, generic dataset (the source) is adapted to a more specialized task (the target) by retraining some or all layers, thereby leveraging learned representations while adjusting to new data characteristics [61, 62].
In seismic segmentation, generalization to new domains is particularly difficult due to variability in acquisition parameters, stratigraphy, frequency content, and noise levels across surveys. This is true for fault delineation, where subtle and discontinuous features are easily obscured by processing artifacts or geologic heterogeneity [63, 47]. Models trained on one domain (such as synthetic datasets like FaultSeg3D [1]) often fail to transfer effectively to real data from different basins. To mitigate such domain shift, researchers have explored domain adaptation techniques including feature-space alignment [64], adversarial learning [65], and seismic-style transfer [66]. Fault-specific approaches also include synthetic fault injection [67] and contrastive learning tailored to structural continuity [20]. Transfer learning pipelines typically rely on pretrained encoders (either from natural images [68, 69] or large-scale seismic simulations with domain-specific decoders [47, 1]) followed by fine-tuning on limited labeled sections. These strategies must navigate a tradeoff between plasticity and stability: aggressive weight updates are prone to inducing catastrophic forgetting [23], while conservative tuning may fail to capture domain-specific features. As a result, generalization in seismic fault delineation requires careful calibration of both model architecture and adaptation strategy to handle the subtle, spatially sparse structures across diverse geological, acquisition, and imaging settings.
II-D Inferential Behavior
In semantic segmentation, inference entails mapping learned feature representations to discrete pixel‐level predictions [28]. The fidelity of this mapping depends on the architecture’s ability to aggregate context and fuse features: models with narrow receptive fields may accurately localize edges but miss global structure [31], while those with extensive context capture coherent regions at the expense of sharp boundaries [29]. During inference, the bias–variance trade‐off emerges as a tension between boundary precision and region consistency, sometimes mitigated by post‐processing, e.g., Conditional Random Fields [70] or edge‐aware refinement modules [71].
In seismic fault delineation, these inferential tendencies are amplified by the thin, curvilinear nature of faults and the high noise intrinsic to seismic volumes [63]. U-Net’s symmetric encoder–decoder and skip connections excel at preserving local detail, producing crisp fault traces when the signal‐to‐noise ratio is high [31], but its reliance on local convolutions and fixed strides can fragment continuous faults under heteroskedastic noise [47]. DeepLab’s atrous convolutions and Atrous Spatial Pyramid Pooling (ASPP) module gather multi‐scale context, yielding smoother, globally coherent fault masks [29]; yet the dilation patterns can inadvertently merge adjacent non‐fault discontinuities, introducing false positives along stratigraphic horizons [29]. SegFormer’s transformer‐based encoder captures long‐range dependencies and adapts to complex fault geometries, enhancing continuity across sections [39]—though its patch‐based attention can produce coarser boundaries if patch size is not carefully chosen [39]. These model‐specific inferential biasesdrive distinct behaviors in fault delineation and inform the choice of architectures and post‐processing strategies for robust seismic interpretation.
II-E Performance Evaluation
Segmentation evaluation seeks to quantify how closely a predicted mask $P$ approximates the ground-truth mask $G$ as subsets of the image domain. Evaluation metrics are categorized in the following two ways: (i) region based measures, which treat masks as sets and quantify volumetric overlap, and (ii) distance-based measures, which treat masks as shapes and quantify geometric deviations along their contours.
Region-based metrics include the Jaccard index (also known as intersection-over-union):
$\mathrm{Jaccard}(P,G)=\frac{\lvert P\cap G\rvert}{\lvert P\cup G\rvert},$ (5)
and the Dice coefficient:
$\mathrm{Dice}(P,G)\;=\;\frac{2\,\lvert P\cap G\rvert}{\lvert P\rvert+\lvert G\rvert},$ (6)
which captures the proportion of shared pixels but may understate errors on thin or tortuous structures.
Distance-based metrics such as Hausdorff offer tolerance against spatial shifts, but they are sensitive to noisy outliers [72]. To mitigate this issue, the modified Hausdorff was proposed by Dubuisson and Jain [73] that considers the average distance to the nearest neighbor, rather than the maximum distance as in the traditional formulation. In [74], the authors utilize the Bidirectional Chamfer Distance (BCD) metric, which measures the minimum distance from each pixel in both directions. Although region based metrics such as the Dice coefficient are useful for assessing shape quality, from a geophysicist’s perspective, evaluation should measure the structural similarity between lines rather than discrete points. The modified Hausdorff distance $D_{h}$ and the BCD $D_{c}$ between two non-empty sets $P$ and $G$ are defined respectively as:
$D_{h}\bigl{(}P,G\bigr{)}$ $=\max\Biggl{(}\frac{1}{N_{p}}\sum_{p\in P}\min_{g\in G}\|p-g\|,$
$\hskip 41.00012pt\frac{1}{N_{g}}\sum_{g\in G}\min_{p\in P}\|g-p\|\Biggr{)}$ (7)
$D_{c}(P,G)$ $=\frac{1}{N_{p}}\sum_{p\in G}\min_{g\in G}\|p-g\|^{2}+\frac{1}{N_{g}}\sum_{g\in G}\min_{p\in P}\|g-p\|^{2}$ (8)
where:
[…]
  • $\min_{g\in G}\|p-g\|$ is the distance from point $p\in P$ to its nearest neighbor in $G$,
  • $\sum_{p\in P}$ and $\sum_{g\in G}$ are summations over all points in sets $P$ and $G$, respectively,
  • ${N_{p}}$ and ${N_{g}}$ are the number of elements in the sets $P$ and $G$, respectively.
To illustrate the complementary nature of these metrics, we present a toy example in Fig. 4. The Dice coefficient measures spatial overlap by comparing each predicted pixel against the ground truth, yielding a single overlap score. In contrast, both the bidirectional Chamfer distance (BCD) and the modified Hausdorff distance assess boundary discrepancies: BCD sums the forward and reverse distances, whereas the modified Hausdorff takes the larger of the two. Because BCD and the modified Hausdorff yield nearly identical trends, our figures plot either one of them alongside the Dice coefficient.
Figure 4: Differences between the considered metrics.
III Fault Delineation Datasets
Among all $74$ datasets used for fault delineation [75], only $4$ field datasets (LANDMASS [76, 9, 77, 78, 79, 80, 81, 82], GSB [45, 83, 84], Thebe [84, 85], CRACKS [86]) and $4$ synthetic datasets (FaultSeg3D [1], Bi’s 3D synthetic [87], Wu’s 2D SR [88], Pochet’s 2D synthetic [89]) open-sourced both seismic data and labels. The low ratio of open-source labeled field data hinders the creation of benchmarks for training and evaluation of models.
Besides the lack of public availability, the different characteristics of the datasets pose challenges to the development of generalizable ML models. Specifically, LANDMASS contains image-level fault labels that cannot be used to numerically evaluate the delineation of pixel-wise faults. GSB contains pixel-wise fault labels annotated on only $5$ crossline sections, which limits the evaluation scalability on large test data. Additionally, it is challenging to achieve generalization using only a small number of labels for finetuning [50]. In contrast, Thebe provides a large amount of pixel-level geophysicist labels across $1803$ crossline sections. CRACKS provides fault labels of varying quality collected from a group of interpreters, including a geophysicist expert, across $400$ inline sections.
Among the four publicly available datasets, we select Thebe and CRACKS considering the model development and evaluation at the pixel level. All the seismic sections in Pochet’s 2D synthetic contain only one straight fault crossing the entire section, presenting less diversity in angle and density compared to the faults in FaultSeg3D. Both Bi’s 3D synthetic and Wu’s 2D SR [88] originate from FaultSeg3D using the same synthesizing workflow. Thus, we use FaultSeg3D as a reference synthetic dataset with a diverse set of faults. Below we describe and compare the three datasets used for our benchmarking study in detail.
FaultSeg3D is a 3D synthetic dataset with $220$ volumes each with dimensions of $128\times 128\times 128$ [1] . In order to better approximate realistic conditions, the authors added background noise estimated from real seismic volumes. The sampling rate and the frequency of the synthetic data vary across the volume to improve the diversity of the data. An example volume is shown in Fig. 5(a).
CRACKS is an open-source dataset with diverse faults delineated across $400$ inline sections of the Netherlands F3 Block [86, 90]. The authors in [10] open-sourced a fully-annotated 3D geological volume of the Netherlands F3 Block for training different models and comparing the performance with objective metrics. Thus, this volume is one of the most extensively studied geographical zones for developing ML-assisted seismic interpretation frameworks [2, 48, 50, 22, 91, 49, 20, 52, 12, 13, 92, 93, 94, 95, 53, 96, 14, 97, 98, 82, 81, 99, 100, 101, 10, 44, 45, 102, 46, 103, 104, 105, 106]. The diverse fault features in the F3 block, including major versus minor faults and varying orientations, make it an excellent seismic dataset to train and evaluate fault delineation models [107, 108]. However, the annotations in [10] do not provide pixel-wise fault labels. Thus, CRACKS open-sources fully hand-labeled fault annotations by a group of $32$ interpreters with varying degrees of expertise and a domain expert geophysicist. This dataset not only establishes a standardized benchmark for objective comparison but also can be used to investigate the impact of multiple sets of labels with varying quality. Three sets of fault annotations are used to investigate the impact of training labels with varying quality, including expert labels and two more sets of lower quality labels from two other annotators. Fig. 5(b) shows an example inline section with expert fault labels from CRACKS. CRACKS and FaultSeg3D share geological similarity in the seismic sections in addition to the similar label density.
Thebe is taken from a seismic survey called Thebe Gas Field in the Exmouth Plateau of the Carnarvan Basin on the NW shelf of Australia [11, 84]. The dataset contains $1803$ labeled crossline sections of size $1737\times 3174$, making it the largest publicly-available field dataset. The seismic intensity of this dataset exhibits a low variation/standard deviation, which can be observed from the low contrast in Fig. 5(c).
(a) FaultSeg3D
(b) Cracks
(c) Thebe
Figure 5: Visuals from three datasets: (5(a)) synthesized faults in FaultSeg3D, (5(b)) expert labels in CRACKS, and (5(c)) expert labels in Thebe.
IV Experimental Setup
The workflow of ML-assisted fault delineation involves multiple choices including ($i$) the decision to fine-tune or not, ($ii$) the selection of the datasets for pre-training and fine-tuning, ($iii$) the selection of different models, ($iv$) the development of pre-processing and post-processing strategies, and ($v$) the standardization of the evaluation protocols. We depict our systematic benchmarking pipeline in Fig. 6, which we use to holistically compare the impact of every choice and the various combinations of such choices. We provide details on each component of our pipeline in this section.
Figure 6: The block diagram of our experimental setup.
IV-A Data Preparation
IV-A1 Standardizing Fault Annotations using Morphological Operations
There exist significant differences between the thickness of fault annotations in CRACKS, Thebe, and FaultSeg3D. We show three labeled patches of the same size from Thebe, CRACKS, and FaultSeg3D in Fig. 7(a). The faults manually delineated by interpreters in Thebe and CRACKS are thicker than the synthesized faults in FaultSeg3D. The inconsistent thickness of fault annotations often influences model optimization, which typically uses pixel-based loss functions. However, existing workflows neglect this fault annotation inconsistency [109, 88, 84, 85]. The synthesized faults in FaultSeg3D precisely delineate the thin fractures in the seismic volumes. We therefore standardize the fault thickness across datasets by processing the manually delineated faults in Thebe and CRACKS to have the same thickness of the faults in FaultSeg3D. Specifically, we apply a simple sequence of morphological operations: we first skeletonize the raw fault annotations and then apply dilation with a rank-3 structural element to uniformize fault thickness and close any small gaps, as shown in Fig. 7(b). This process is denoted by the Preprocess block in Fig. 6. We use standard Python libraries scikit-image and scipy for these morphological operations.
IV-A2 Training and Test Splits
To ensure meaningful evaluations, we adopt a consistent splitting strategy across the considered datasets. Specifically, we maximize diversity in the test set while ensuring no overlap with the training data. For each dataset, we select spatially distinct subsets to prevent redundancies and simulate deployment conditions on previously unseen segments. In CRACKS, which consists of 400 contiguous inlines, we designate the first 30 and last 30 sections as the test set, totaling 60 sections. The remaining 340 central sections are used for training. This split captures geological variability across the volume while maintaining spatial separation between training and test data. For Thebe, we use 400 sections for training and reserve 100 sections from other parts of the volume for testing, following a similar diversity-maximizing approach. For FaultSeg3D, we follow the established setup from [1], using 200 synthetic volumes for training and 20 distinct volumes for testing.
(a) Fault labels before morphological operations
(b) Fault labels after morphological operations
Figure 7: Example $128\times 128$ patches of fault annotations with varying thickness in the three datasets. Left: Thebe, middle: CRACKS, right: FaultSeg3D
IV-B Model setup
We consider eight segmentation models commonly used in the seismic literature [84, 88, 5, 110] in our fault delineation experiments, using the following nomenclature:
  • deeplab: Deeplab [111] architecture with a Resnet50 [68] backbone
  • deeplab-m:Deeplab architecture with a Mobilenet [112] backbone
  • hed: Hollistically-nested edge detection model [113]
  • rcf: Richer convolutional features [114]
  • unet: Unet [31] architecture with a Resnet50 backbone
  • unet++: Unet++ [115] architecture with a Resnet50 backbone
  • unet-b: Original Unet architecture as presented in [31]
  • segformer: Transformer-based model introduced in [39]
For each of these models, we perform all pairwise combinations of pretraining-finetuning settings between the $5$ sets of labeled training data in the top left of Fig. 6. For example, a model is first pretrained on CRACKS-expert and then finetuned on all $5$ datasets, resulting in $25$ sets of weights for each considered model. Each of these model weights is then evaluated on CRACKS, Thebe, and FaultSeg3D. To complete our baselines, we also use four ImageNet-pretrained models and finetune them on our seismic datasets.
IV-C Evaluation Setup
IV-C1 Post-processing for Model Predictions
Common practices of evaluating deep fault delineation networks involve two steps [50, 8, 85, 84]: $(i)$ thresholding the network predictions to obtain binary outputs, and $(ii)$ comparing the binary outcomes with the fault test labels using a metric. The test labels are processed with the same morphological operations as the training faults in order to achieve consistent fault thickness at the input and output of a model. Consequently, the thresholding in step $(i)$ needs to be adaptive to the model and the data accordingly, followed by the same morphological operations for numerical evaluation. We compute the optimal threshold for the dataset using the Optimal Dataset Scale (ODS) metric [109]. For a model, its optimal threshold is computed using the training set, which is then applied to binarize the predictions on the test data, followed by the same morphological operation for evaluation.
IV-C2 Performance Metrics
As mentioned in Section II-E, pixel-based and distance-based metrics capture different aspects of prediction quality, and each can fail to fully reflect the structural accuracy of the predicted faults, making the evaluation of fault delineation methods an inherently challenging task. Our benchmark study thus adopts a holistic approach to assess prediction quality by considering a combination of multiple metrics alongside subjective inspection. In this study we choose to report one pixel-based metric, Dice Coefficient (defined in Eq. (6)) and two distance-based metrics: BCD and the modified Hausdorff (defined in Eq. 8, and Eq. 7, respectively), whose difference we showcased in Fig. 4. These metrics have not only been extensively used in the seismic literature for model performance assessment [1, 25, 73], but allow for two different evaluation axes: pixel overlap-based and structure-based.
V Results
Benchmarking results are analyzed across three different thematic axes: (1) generalizability and transferability, (2) training dynamics, and (3) metric evaluation.
Figure 8: Visual encoding to represent the different models pretrained and finetuned on different datasets. Border color signifies the pretraining data source and the fill color signifies the fine-tuning dataset. Models are represented using various shapes.
V-A Generalization and Transferability
As mentioned in Section I, models are often used to process data from new surveys that can differ from their original training sources. As such, generalization and transferability are critical for reliable deployment and to reduce the labeling overhead. Understanding how various training regimes perform across domains is key to developing scalable workflows.
In this subsection, the generalization of models trained under different pretraining-finetuning regimes is analyzed across CRACKS, FaultSeg3D, and Thebe. Unless stated otherwise, we structure our scatter plots using the convention depicted in Fig. 8. The shape of a given point encodes the model used, the border color corresponds to the dataset used for pretraining the model, and the fill color corresponds to the datasets used for finetuning the model. All of our reported figures and plots correspond to models being evaluated in a held-out test partition of one of these 3 datasets.
V-A1 Dataset Alignment and Transfer Trends
Figs. 910 and 11 provide a macroscopic view of the model behaviors across the different training setups. On CRACKS (Fig. 9), the top-performing configurations are those pretrained on FaultSeg3D data and fine-tuned on CRACKS. The observation indicates strong geophysical commonalities between FaultSeg3D and CRACKS data, and supports the utility of FaultSeg3D data as a viable pretraining source when real annotations are limited. The fill-color distribution in Fig. 9 also establishes a hierarchy of effective fine-tuning datasets for CRACKS: CRACKS ¿ FaultSeg3D ¿ Thebe. The poor finetuning performance on Thebe indicates that it is more distributionally distant from CRACKS than FaultSeg3D data.
Furthermore, when tested on FaultSeg3D data (Fig. 10), models pretrained on CRACKS again outperform those trained from scratch, indicating that the aforementioned alignment is reciprocal. However, for Thebe (Fig. 11), top-performing models are those trained from scratch on Thebe itself. Transferring from either FaultSeg3D or CRACKS results in performance degradation, suggesting that Thebe resides in a distinct feature space.
Figure 9: Test on CRACKS dataset
Figure 10: Test on FaultSeg3D dataset
Figure 11: Test on Thebe dataset
V-A2 Domain Shift and Joint Training
Domain shift is a pressing challenge in seismic ML, particularly when deploying models across surveys with different geological properties. When overlooked, this can lead to brittle models that perform well on training data but fail on new volumes, or models that catastrophically forget features from their original data after finetuning. An illustration of this catastrophic forgetting occurs when models pretrained on CRACKS are fine-tuned on Thebe: despite reasonable performance on Thebe, these models experience dramatic performance degradation when re-evaluated on their original domain. In the case of unet++, for instance, the Dice score on CRACKS drops from 0.34 to 0.12, while the BCD increases twofold, indicating a complete erasure of useful representations, and clear case of catastrophic forgetting. A statistical analysis shows that Thebe has a significantly lower standard deviation (0.124) compared to CRACKS (1.149) and FaultSeg3D (1.052). The contrast and intensity variations in Thebe are lower, while CRACKS and FaultSeg3D have diverse intensity distributions, as shown in Fig. 5. An explanation for this catastrophic forgetting phenomenon is that models trained on Thebe learn from a constrained input range and struggle when tested on datasets with richer intensity distributions. These results suggest that other normalization techniques before training could help reduce these distributional mismatches.
As an additional baseline, experiments on joint training on all possible combinations of datasets in a single training round are conducted using the unet model with a ResNet50 backbone. The results for all unet experiments, including these joint training configurations, are shown in Table I. The results further showcase that features learned from CRACKS and FaultSeg3D data align together, but these learned features are not easily transferable to Thebe. On the other hand, Thebe acts as a regularizer, hurting performance on seen datasets but providing modest generalization to unseen ones.
TABLE I: Performance Metrics (DICE, BCD, Hausdorff) for different training schemes on Unet. All values are rounded to three decimals. Schemes are ranked as best, second best or worst based on satisfying at least two of the three metric criteria, to account for cases where metrics disagree. (blue for highest, pink for second, and gray for the worst.)
Training Configuration Test on FaultSeg3D Test on Cracks Test on Thebe
DICE BCD Hausdorff DICE BCD Hausdorff DICE BCD Hausdorff
Individual Training
Thebe 0.111$\pm$0.036 18.767$\pm$18.819 16.095$\pm$18.858 0.020$\pm$0.008 69.043$\pm$15.360 49.012$\pm$15.350 0.182$\pm$0.019 52.054$\pm$7.696 31.615$\pm$6.243
FaultSeg3D 0.416$\pm$0.059 9.563$\pm$9.446 7.191$\pm$7.769 0.160$\pm$0.046 27.016$\pm$7.382 20.638$\pm$6.683 0.042$\pm$0.008 232.721$\pm$28.861 189.939$\pm$27.003
Cracks 0.012$\pm$0.054 76.904$\pm$28.473 57.798$\pm$18.930 0.333$\pm$0.067 11.330$\pm$6.431 7.341$\pm$4.732 0.001$\pm$0.003 1377.945$\pm$143.872 1022.510$\pm$114.634
Combined Training
FaultSeg3D + Cracks 0.662$\pm$0.190 6.143$\pm$20.421 5.282$\pm$18.295 0.333$\pm$0.048 15.436$\pm$5.254 11.597$\pm$4.836 0.001$\pm$0.001 659.708$\pm$74.985 430.887$\pm$64.880
FaultSeg3D + Thebe 0.060$\pm$0.031 29.834$\pm$20.870 20.151$\pm$19.805 0.167$\pm$0.033 29.717$\pm$8.512 26.136$\pm$8.187 0.211$\pm$0.024 52.952$\pm$10.802 43.348$\pm$10.516
Cracks + Thebe 0.038$\pm$0.3375 34.183$\pm$28.569 22.751$\pm$23.675 0.32$\pm$0.067 13.797$\pm$6.2072 8.2964$\pm$4.64058 0.112$\pm$0.03 112.354$\pm$38.606 60.198$\pm$40.66
All 0.337$\pm$0.035 28.569$\pm$13.036 23.674$\pm$11.087 0.306$\pm$0.0743 17.7453$\pm$21.851 12.005$\pm$19.724 0.144$\pm$0.029 75.877$\pm$37.191 59.039$\pm$19.8
Fine-Tuning
Cracks $\rightarrow$ FaultSeg3D 0.667$\pm$0.171 4.994$\pm$16.049 3.821$\pm$10.475 0.028$\pm$0.009 73.433$\pm$13.961 45.378$\pm$13.263 0.027$\pm$0.003 165.315$\pm$10.283 99.996$\pm$7.698
Cracks $\rightarrow$ Thebe 0.099$\pm$0.036 21.959$\pm$13.845 15.771$\pm$10.753 0.159$\pm$0.028 26.070$\pm$7.107 20.929$\pm$6.552 0.181$\pm$0.014 53.808$\pm$10.361 42.169$\pm$12.450
FaultSeg3D $\rightarrow$ Cracks 0.082$\pm$0.080 136.293$\pm$151.136 134.095$\pm$151.213 0.375$\pm$0.059 8.806$\pm$3.728 5.577$\pm$2.71 0.019$\pm$0.005 213.695$\pm$20.154 147.213$\pm$12.756
FaultSeg3D $\rightarrow$ Thebe 0.075$\pm$0.082 93.488$\pm$66.848 84.420$\pm$61.439 0.135$\pm$0.023 28.701$\pm$7.022 23.368$\pm$7.111 0.158$\pm$0.014 78.683$\pm$35.415 51.649$\pm$30.973
Thebe $\rightarrow$ Cracks 0.019$\pm$0.027 56.145$\pm$34.593 42.899$\pm$34.785 0.317$\pm$0.047 10.896$\pm$3.127 6.656$\pm$2.500 0.019$\pm$0.005 161.131$\pm$14.429 88.344$\pm$9.217
Thebe $\rightarrow$ FaultSeg3D 0.526$\pm$0.203 9.319$\pm$22.281 8.182$\pm$21.079 0.053$\pm$0.012 38.141$\pm$7.286 34.319$\pm$7.233 0.037$\pm$0.003 134.522$\pm$7.354 122.936$\pm$7.078
V-A3 Model Capacity and Transferability
It is generally established in the literature that pretraining on a large dataset can boost the performance of a model even in self-supervised settings [5]. However, in seismic applications, the benefits of pretraining are critically sensitive to the distributional similarity between the source and target domains—a phenomenon that reflects the strong coupling between geologic variability and model transferability. As shown in our experiments, models pretrained on FaultSeg3D or CRACKS and fine-tuned on Thebe underperform compared to models trained from scratch or initialized from ImageNet weights (Fig. 12). In such distributionally mismatched cases, the standard pretaining-finetuning strategy may hinder rather than help performance.
A factor that also plays a role in modulating generalization is model capacity. Fig. 12 and Fig. 13 show that large models like segformer benefit from pretraining on CRACKS, while smaller models (e.g., hed, deeplab-m) generalize better when trained from scratch. The observation suggests that both data alignment and model capacity affect transfer effectiveness.
Furthermore, models respond differently to fine-tuning. When pretraining on FaultSeg3D and fine-tuning on Thebe, models degrade in CRACKS performance. Fig. 14) shows that Thebe induces domain shifts that are hard to unlearn. Conversely, even though the finetuning dataset is distributionally closer to the target, Thebe-pretrained segformer and deeplab also degrade on CRACKS after finetuning on FaultSeg3D as shown in Fig.  14. The asymmetric behavior highlights the difficulty of finding universally robust pretraining strategies.
Figure 12: Individual Models tested on Thebe. Models with blue edges are pretrained on Thebe without finetuning. While other models are pretrained on different datasets. We show that pretraining on another faults dataset is not beneficial compared to using Imagenet weights or training from randomly initialized models.
Figure 13: Individual models tested on Thebe. (a) Models pre-trained on the FaultSeg3D data, and fine-tuned on different data. (b) Models pre-trained on the CRACKS data, and fine-tuned on different data.
Figure 14: Individual models tested on CRACKS. (a) Models pre-trained on the FaultSeg3D data, and fine-tuned on different data. (b) Models pre-trained on the Thebe data, and fine-tuned on different data.
V-B Training Dynamics and Inferential Behavior
In this section we analyze the impact of different training dynamics and model architectures on performance and inferential behavior.
V-B1 Window Size and Loss Function
We evaluate the unet model with a ResNet50 backbone using Dice and Binary Cross-Entropy (BCE) losses, as well as across different window sizes: 96, 128, 256 and 512, or up to the size allowed by the original sections in each dataset. The results for these experiments are shown in Fig. 15, where bigger markers correspond to bigger patch sizes. We can see that when using Dice as a loss, models benefit from having access to larger window sizes, while the same does not hold when using the BCE loss. This discrepancy is due to the class imbalance nature of the fault delineation task: BCE tends to work best when foreground and background are balanced but fails otherwise, and Dice is generally designed for class-imbalanced scenarios [116]. In this context, using smaller patches is equivalent to ’zooming into’ the small fault regions and therefore reducing the class imbalance, thus improving BCE performance. Overall, this further validates our choice of using the Dice loss on our benchmarking experiments.
Figure 15: Behavior of DICE and BCE losses under different window sizes
V-B2 Model Nuances in Fault Delineation
We also qualitatively observe that each of the evaluated models presents different nuances in the structure of their fault predictions, irrespective of the pretraining or finetuning strategy used, many of which can be observed in Figs. 19 and 20. For example, deeplab tends to produce irregular, stair-like faults, while segformer produces thicker, blob-like faults. unet architectures in general tend to produce thinner faults, with unet++ generating more fragmented ones. These architectural signatures are consistent across training setups and highlight the influence model design choices inherently have in shaping the morphology of predicted faults, which is an important consideration when selecting models for downstream tasks or when interpreting evaluation results beyond numerical metrics.
V-C Metric Robustness and Observability
Due to the high correlation among adjacent sections in a seismic volume, deep learning models tend to generate consistent patterns that vary minimally between neighboring sections. Evaluation metrics respond differently to these subtle variations; some metrics heavily penalize these deviations, while others are more tolerant. The effects of structural variations in fault predictions on the evaluation metrics are investigated.
V-C1 Sensitivity to Visual Structure
Distance-based metrics are generally more tolerant to the structure of the predicted fault. This behavior is illustrated in Fig.  LABEL:fig:dicevsbcd. Where Fig. LABEL:fig:dicevsbcd_a represents the ground truth, while LABEL:fig:dicevsbcd_b and LABEL:fig:dicevsbcd_c show the predictions of two different models. Although the prediction in Fig. LABEL:fig:dicevsbcd_b appears structurally closer to the ground truth, it receives significantly worse BCD and Hausdorff scores compared to the prediction in Fig. LABEL:fig:dicevsbcd_c. Since distance-based metrics do not heavily weigh continuity, the inclusion of a few extra pixels around the fault can improve the BCD score even if those pixels lack proper structural alignment. Notably, these patterns are not anomalies, different models often generate such consistent outputs. Consequently, numerical metrics can be misleading and may display discrepancies between one another.
Figure 16: Example of the structure tolerance in distance-based metrics. (a) shows the ground truth fault annotations. (b) and (c) show predictions from two different models. While (b) appears structurally closer to the ground truth, it receives a significantly worse BCD and Hausdorff but a better Dice score.
V-C2 Fault Sparsity
Another issue with distance-based metrics is that they are designed to measure the quality of a single continuous object in the image, whereas faults can consist of multiple sparse objects. This makes distance-based metrics very sensitive to both the number of faults present in the ground truth and their sparsity. We showcase this issue in Fig.  17, where we consider two cases: one with a few faults and another with many faults. Noise is added to both images, and each is compared against its original version. The case with fewer faults contains many outliers compared to the case with many faults, since the added noise often lies far from any existing fault. As demonstrated in Fig.  4, because BCD is bidirectional, it accounts for each added noise pixel by searching for its closest existing fault. Therefore, having fewer faults results in a much worse BCD score. DICE, while also penalizing noise, is more stable across different sparsity levels.
Figure 17: Sensitivity of distance-based metrics to fault density: Sections with many (a) vs. few faults (b) are evaluated under identical noise. Despite equal noise levels, (b) is penalized more by distance-based metrics, while Dice remains stable.
V-C3 Contradictory Scores and Human Judgement
Although pixel-based metrics are more stable, they remain sensitive to slight pixel shifts. In Fig.  18, we show examples corresponding to two different models. The prediction in Fig. 18(c) looks visually closer to the ground truth, but receives a worse Dice score (0.1325 vs. 0.13421) and significantly better BCD and Hausdorff (40.782 vs. 115.436 and 29.0 vs 72.5). Since the faults in both predictions are poorly structured and spatially misaligned, the Dice coefficient (being overlap-based) penalizes them similarly. In contrast, distance-based metrics such as BCD and Hausdorff distances are more sensitive to the spatial coherence of the predicted faults, hence, they are more tolerant to the structure of the faults. These contradictions imply that some metrics may conflict with visual intuition or downstream utility, and point to the need for context-aware metric selection frameworks.
(a) Ground Truth
(b) BCD:115.43, Hausdorff: 72.5, Dice: 0.1342
(c) BCD:40.782, Hausdorff: 29.0, Dice: 0.1325
Figure 18: (a) Ground truth fault annotations. (b) and (c) display predictions from two different models. Although the prediction in (c) appears visually more aligned with the ground truth than (b), it receives a slightly worse Dice score but substantially better BCD and Hausdorff distance
Figure 19: Predictions of the models pretrained on FaultSeg3D tested CRACKS
Figure 20: Predictions of the models pretrained on CRACKS tested Thebe
TABLE II: Key observations
Topic Key Findings
Distributional shift among seismic data The intensity standard deviation of the synthetic FaultSeg3D data is similar to that of CRACKS, while both are very different from Thebe’s. This can be attributed to the discrepancy of seismic features across datasets.
Relationship between CRACKS and FaultSeg3D Both CRACKS and FaultSeg3D data benefit from pretraining on the other, outperforming models trained from scratch in either.
Thebe vs. others Thebe does not benefit significantly from pretraining on other data; training from scratch performs best due to distributional shifts.
Model size relationship with pretraining Larger models like segformer and unet (ResNet50) perform well when pretrained on other datasets and finetuned on FaultSeg3D. Smaller models like rcf and hed degrade in performance with pretraining, indicating a lack of transfer capacity.
Fault density FaultSeg3D and CRACKS have dense faults; Thebe faults are sparse, affecting model prediction density.
Metric structural biases Dice penalizes misshaped faults more, while Hausdorff/BCD may still give high scores due to proximity.
Metric sparsity biases Fewer faults lead to harsher penalty in distance metrics; dense faults often score better.
deeplab behavior Produces jagged or stair-like faults.
segformer behavior Tends to generate thick, blob-like faults.
unet/unet++ behavior unet creates thin faults; unet++ tends to produce fragmented ones.
hed/rcf behavior Less adaptable to Thebe due to fault density mismatch; outputs noisy, distorted shapes.
Joint CRACKS-FaultSeg3D Combining CRACKS and FaultSeg3D data leads to synergistic features and better results.
Joint training with Thebe Adding Thebe acts as a regularizer: performance drops on original domains but improves generalization.
VI Conclusion
In this work, we present the first large-scale benchmarking study of machine learning models for seismic fault delineation under domain shift. By systematically training and evaluating over 200 models across three diverse seismic datasets, we provide a comprehensive analysis of generalization, transferability, and evaluation challenges in seismic segmentation. Our results reveal the fragility of fine-tuning strategies when faced with shifts in data distribution, the emergence of catastrophic forgetting when moving across domains, and the sensitivity of model behavior to architecture and pretraining alignment. We also highlight the limitations of standard evaluation metrics, which often fail to capture geological plausibility, and we advocate for the development of more nuanced frameworks that account for expert ambiguity and structural realism. Collectively, these findings establish a robust empirical foundation for understanding the tradeoffs in current seismic ML workflows and offer practical guidance to researchers and practitioners seeking to build models that are not only accurate but also transferable, interpretable, and aligned with real-world exploration needs.
Acknowledgment
This work is supported by ML4Seismic Industry Partners at the Georgia Institute of Technology.