### v4

# This work is supported by the ML4Seismic Industry Partners at Georgia Tech

Abstract
Machine learning has taken a critical role in seismic interpretation workflows, especially in fault delineation tasks. However, despite the recent proliferation of pretrained models and synthetic datasets, the field still lacks a systematic understanding of the generalizability limits of these models across seismic data representing diverse geologic, acquisition, and processing settings. In practice, distributional shifts between surveys, limitations in fine-tuning strategies, and inconsistent evaluation protocols remain major obstacles to deploying reliable models in real exploration settings. In this paper, we present the first large-scale benchmarking study explicitly designed to evaluate domain shift strategies for seismic fault delineation. Our benchmark spans over 200 experimental setups combining eight architectures, three datasets (FaultSeg3D, CRACKS, Thebe), and multiple training regimes (individual training, fine-tuning, and joint training). We evaluate performance using three complementary metrics (Dice, Hausdorff Distance, and Bidirectional Chamfer Distance) and analyze both segmentation accuracy and structural fidelity. Our results show that fine-tuning across dissimilar domains can reduce source-domain Dice by up to 75%, demonstrating severe catastrophic forgetting, whereas larger models such as Segformer tend to be more robust to adaptation than smaller architectures. We also find that domain adaptation methods outperform fine-tuning under large distributional gaps but underperform when domains are closely aligned. Finally, we complement conventional metrics with an analysis of fault-network descriptors (length, curvature, sinuosity, segmentation density), revealing the nuances in the interplay between architectural choices and data properties. Overall, this benchmark provides a reproducible foundation for evaluating transferability in seismic fault delineation and offers actionable insights for effective deployment of fault delineation models within seismic interpretation workflows.
Index Terms:
Benchmarking, Domain shift, Seismic fault delineation, Seismic interpretation
[…]
=-15pt
I Introduction
Recent advances in deep learning (DL) have brought about a tectonic shift in the seismic interpretation workflow [isah_review_2025, wenxue_overview_2025], mirroring broader trends across other domains like geoscience [yaqoob_geocrack_2024, qaiser_vision_2025], sustainable energy systems [yaqoob_advancing_2025, yaqoob_microcrystalnet_2025, yaqoob_fluidnet-lite_2025] and biomedical applications [logan2022multi, logan2022patient, kokilepersaud2023clinically, quesada2022mtneuro, kokilepersaud2022gradient, iffa2025development, ansari_survey_2025]. DL-assisted approaches are leveraged in different parts of the pipeline, specifically in automated fault detection. Faults are geologically critical features that control fluid flow in Earth’s subsurface, influence reservoir compartmentalization, and pose drilling hazards in hydrocarbon exploration [wu2019faultSeg, chowdhury2025unified]. Beyond the energy sector, faults play a central role in earthquake nucleation and propagation, as well as geohazard and risk assessment in tectonically active regions [benzion2008collective, KANAMORI20031205]. Accurate and scalable fault interpretation is therefore a high-impact task across several geoscience domains. Early work in this area often repurposed semantic segmentation models from computer vision (such as those developed for natural images) to detect discontinuities in seismic sections. However, these generic frameworks struggled with the unique characteristics of seismic data, particularly the thin, curvilinear geometry of faults and the presence of acquisition artifacts and structural noise. As a result, the field has increasingly shifted toward fault delineation, a task-specific variant of segmentation that emphasizes the extraction of coherent fault structures rather than general pixel-wise classification. This shift has prompted the development of domain-specific neural networks, feature encoders, adapted loss functions, and structural priors that better capture the morphological signatures of faults [rs16050922, faultsam]. In general, the growing availability of DL methods has accelerated the adoption of data-driven interpretations that have provided new insights into fault structures within large seismic volumes [rs16050922, faultsam].
Figure 1: Typical DL-assisted seismic interpretation workflow
A DL-assisted seismic interpretation pipeline is illustrated in Fig. 1 (top diagram). In a standard workflow, interpreters at a new site begin by leveraging existing labeled datasets or synthetically generated volumes to train machine learning models. These models capture broad geophysical patterns and serve as a foundational prior, which can then be adapted to the new site through fine-tuning on a smaller set of locally interpreted seismic data. After fine-tuning, the models are evaluated within a consistent framework to ensure that their outputs align with geological expectations and interpretation standards. The validated predictions subsequently guide subsurface decision-making, and the new volume may in turn contribute to the pool of training data for future sites. However, while conceptually straightforward, this workflow embeds multiple assumptions and design choices that present significant challenges in practice, which we highlight at the bottom of Figure 1.
Two central and intertwined challenges in the interpretation pipeline arise from the diversity of data sources (Fig.  1, orange box) and the strategies used to adapt models across them (Fig. 1, green box). Seismic datasets differ widely in their geological characteristics, acquisition parameters, and resolution, among other factors. Synthetic data, while providing a fully specified ground truth, often fails to capture the complexity and variability of field data. Prior datasets from other regions, though more realistic, may still differ significantly from new target volumes. These domain shifts, whether from synthetic-to-real or across real-world basins, can cause pretrained models to generalize poorly to new sites [nasim2020seismic], and eventually provide weak performance for the interpretation task.
Figure 2: Finetuning example from a real dataset (D1) and a synthetic dataset (D2) to a target real dataset
To address the challenge of domain adaptation, researchers have proposed a variety of methods that address the effects of domain differences along the pipeline. Several studies introduce enhancements to model architectures [LI2023105412], dataset construction [alaudah2018structure, alaudah2019machine, wu2019faultSeg, AN2021107219], uncertainty estimation [benkert2022reliable] or explore training strategies aimed at improving data efficiency and transferability [benkert2023samples, mustafa2021man]. A common approach involves pretraining models on synthetic data and subsequently fine-tuning them on real seismic volumes, as in [wu2019faultSeg]. More recent methods have explored using self-supervised learning to learn robust features without large amounts of labeled data [wang2016interactive, cohen2006detection, roberts2001curvature, di2017seismic, shafiq2018towards, kokilepersaud2022volumetric], or adapting large vision foundation models trained on natural images to the seismic domain [faultsam, Quesada_2024_CVPR, quesada2024benchmarking]. While these approaches offer promising results, their effectiveness is often dependent on the compatibility between source and target domains, and they remain vulnerable to issues such as catastrophic forgetting [kirkpatrick2017overcoming] during adaptation. Moreover, the success of these approaches is influenced by several factors such as architecture and training design choices, as well as dataset-specific properties such as seismic section resolution, heterogeneity, and sample size.
Fig. 1 (green box) showcases a simple example of the aforementioned issues. The figure contains two rows, each displaying three seismic sections. The first row presents the output of three segmentation models commonly used in the seismic community, trained under the same settings with the same training data. We can observe that the three predictions differ significantly from each other. In the second row of the box, three segmentation outputs are shown after applying three different fine-tuning strategies. The first output in the second row is obtained by a model trained from scratch on the target data, while the other two are produced by models pretrained on other (larger) datasets and then finetuned on the target data. We further conduct objective evaluation of these outputs and summarize these metrics in Fig. 2. The y-axis in this figure is the average Hausdorff distance between the output of the model and the ground truth, where smaller values signify better predictions. There are four strategies shown. The first one is where the model is trained on the target volume. The second strategy is to use a model that was trained on a real dataset and fine-tune it on the target dataset. The third strategy is to use a model that was trained in a synthetic dataset and then fine-tuned on the target dataset. These three strategies match the three depicted in Fig. 1(green box). A fourth strategy is to use a model that was jointly trained on both synthetic and real data, then fine-tuned on the target data. Although we will discuss these experiments in more detail later in the paper, the figure here shows the large difference in performance across these four strategies. For now, this shows that despite this growing body of work, the field still lacks a systematic understanding of when, how, and why a given strategy outperforms others under domain shift. This paper is the first attempt, to our knowledge, within the community to provide answers and guidelines for domain shift strategies in seismic interpretation.
The third critical complication in the pipeline arises in evaluation, as depicted in Fig. 1 (purple box). Fault labels are derived from expert interpretation and are often subjective [alcalde2017impact, guillon2020ground], particularly in complex or ambiguous structural settings. This subjectivity leads to inconsistencies in ground truth across datasets or among annotators, and complicates objective model evaluation. Furthermore, a given metric may penalize correct but unannotated predictions, while failing to reflect geological plausibility or structural consistency [sarajarvi2020robust, rs16050922]. To address this, our study provides an in-depth analysis of commonly used evaluation metrics (such as Dice coefficient, Chamfer and Hausdorff distances) in the context of fault interpretation. This analysis highlights the need for a more holistic evaluation approach that integrates quantitative measures with geological interpretability.
This paper presents a large-scale benchmarking study regarding the performance variability of finetuning strategies for fault delineation models across training regimes, data characteristics, and architectural choices. Rather than introducing a new model, this work establishes a systematic experimental framework: a methodological foundation for reproducible and cross-domain evaluation of deep learning architectures in seismic fault interpretation. Our benchmarked fault delineation models include a combination of eight deep learning architectures, trained and fine-tuned on three distinct datasets, including both synthetic and real volumes. Our experiments cover a broad spectrum of training strategies, including pretraining on a single fault dataset, pre-training on large-scale natural image dataset such as ImageNet [deng2009imagenet], using randomly initialized models and jointly training on multiple fault datasets. Beyond benchmarking, we also use targeted case studies (through domain adaptation and continual learning setups) to diagnose model behavior under catastrophic forgetting and domain shift, demonstrating how the benchmark can serve as a tool for both evaluation and insight.
While prior works have examined transfer learning and domain adaptation in geoscience [sun2016deep, tsai2018learning, trinidad2023seismic], these efforts have typically been limited in scope: focusing on single architectures or isolated dataset pairs. By contrast, our study provides the first large-scale, systematic benchmark that unifies evaluation across multiple model families, datasets, and training paradigms. This effort parallels the role of domain-shift benchmarks in fields like medical imaging [guan_domain_2022], where standardized comparisons have proven essential for understanding generalization and reproducibility across domains.
Our contributions are as follows:
  • We present the first large-scale benchmarking study focused on fault delineation under domain shift, spanning over 200 exhaustive combinations of experimental setups across 8 DL architectures, 5 pretraining-finetuning datasets and 3 evaluation ones, as well as several training configurations.
  • We systematically evaluate pretraining, fine-tuning, and joint training strategies, exposing failure cases such as catastrophic forgetting and dataset-dependent brittleness.
  • We analyze the effect of architectural scale and design on transferability, highlighting that larger models often adapt more effectively to finetuning, and smaller models benefit from domain adaptation strategies under large domain shift.
  • We compare evaluation metrics and highlight their limitations, advocating for evaluation practices that incorporate structural and geological plausibility. We also introduce a novel analysis based on fault characteristic metrics, which sheds new light on the way data properties and model behavior are entangled, and opens up new avenues from which to understand finetuning and domain transfer strategies.
  • We open-source our code and models to provide a foundational reference to scientists and practitioners.
The remainder of this paper is organized as follows. Section II reviews prior work on machine learning for fault delineation in the context of seismic interpretation workflows. Section III describes the three datasets used in our benchmarking study, highlighting their geological and statistical characteristics. Section IV details our experimental setup, including data preparation, model architectures, and evaluation protocols. Section V presents the results of over 200 training configurations, analyzed across generalization and transferability, training dynamics, and metric evaluation. Section VI introduces an additional analysis on geometric and topological metrics to characterize dataset–model interactions. Finally, Section VII concludes the paper by summarizing key findings, outlining practical guidelines, and identifying open challenges for future research. The full cosde and data used for this paper can be found at https://github.com/olivesgatech/large-bench-geo.
II Related Work
The structure of this section mirrors the stages outlined in Fig. 1. Existing literature on machine learning for fault delineation follows a similar modular approach: beginning with data preprocessing and input normalization, proceeding through model training, adaptation strategies and inferential behavior, and culminating in evaluation protocols. To provide both breadth and depth, each subsection first summarizes the general deep learning principles that motivate a given stage, and then narrows the discussion to prior work in seismic interpretation and fault segmentation specifically. In this way, we highlight how established DL techniques intersect with geophysical challenges, and how each component contributes to generalization across seismic domains.
II-A Data Preparation and Preprocessing
In computer vision segmentation pipelines, raw images from natural scenes [long2015fully, chen2017deeplab, kokilepersaud2023focal] and medical scans [ronneberger2015u, isensee2021nnu, prabhushankar2022olives] are first standardized to ensure consistent spatial dimensions and intensity distributions [ioffe2015batch, ulyanov2016instance]. Typical preprocessing steps include resizing or cropping to a fixed height and width, data augmentation, whitening [prabhushankar2017generating], and contrast adjustments to mitigate variability in lighting or sensor settings [shorten2019survey, wang2017effectiveness]. These operations guarantee that each input conforms to the network’s architectural requirements and that learned features are not biased by extraneous intensity fluctuations.
Seismic segmentation extends these practices to volumetric data acquired in formats such as SEG-Y (.sgy), raw binary (.dat), or serialized NumPy arrays (.npy). A common workflow begins by reading trace headers to assemble a 3D volume of dimensions (inline $\times$ xline $\times$ time/depth), and either processing the volume at that level [wu2019faultSeg, rs16050922] or extracting 2D inline or crossline sections for further processing [alaudah2018structure, di2017seismic]. To enlarge the training set and fit GPU memory constraints, each section is tiled with an overlapping sliding window of size $H\times W$ pixels (the effects of different tiling standards are studied and results presented in Section V-B1.).
Since raw seismic amplitudes can span orders of magnitude and contain acquisition artifacts, it is a standard practice to normalize each volume either via min–max scaling or z-score transformation. Fault masks (either interpreter‐drawn or synthetically generated) are saved as binary images (.png or .npy) and cropped identically to their corresponding seismic windows. This systematic preprocessing pipeline (from raw SEG-Y/DAT ingestion, through standardized normalization and windowed slicing, to precisely aligned input–mask pairs) provides a consistent basis for benchmarking and comparing fault delineation models across diverse seismic domains.
II-B Training Dynamics and Setups
Popular network choices designed for natural image segmentation are sensitive to training and design choices like model architecture, input window size, and loss functions. Networks such as U-Net [ronneberger2015u], DeepLab [chen2017deeplab], and SegFormer [xie2021segformer] typically accept fixed‐size $n\times n$ patches (usually $n\in\{128,256,512\}$), balancing the need for sufficient context against GPU memory limits. During each training iteration, these patches are sampled, sometimes at multiple scales to expose the network to varied object sizes and to regularize against overfitting. Segmentation models optimize losses designed to reconcile pixel‐wise accuracy with region‐level coherence. The binary cross-entropy (BCE) loss [cox1958regression] focuses on classifying each pixel correctly, while the Dice loss [milletari2016v] measures the overlap between predicted and ground-truth fault regions, making it well-suited for imbalanced datasets. In practice, many works combine BCE and Dice to take advantage of both pixel fidelity and shape alignment [wazir2022histoseg, zu2024resaceunet].
In seismic fault delineation, similar principles apply but with additional considerations. Input patches are typically larger to capture fault continuity across sections, and training often uses overlapping windows with a stride smaller than the patch size to ensure boundary faults are adequately sampled [wu2019faultSeg, di2019developing, di2019improving, di2019semi]. The aforementioned loss functions, coupled with learning rate schedules and different regularization functions, form the backbone of seismic training setups [dou2022md, benkert2024effective, mustafa2023active]. By carefully tuning window sizes and loss compositions, researchers mitigate class imbalance and preserve structural continuity in fault delineation [mustafa2024visual].
II-C Generalization and Transferability
Generalization refers to a model’s ability to maintain performance when presented with new, unseen data that (ideally) follows the same distribution as the training set [goodfellow2016deep, benkert2022example, mustafa2021joint]. In segmentation, this means accurately delineating objects or regions under variations in lighting, scale, or background texture. However, a common obstacle to generalization is domain shift, which occurs when the statistical properties of training and test data differ (such as changes in camera sensors or scene composition) often leading to degraded performance [torralba2011unbiased, Temel2017_NIPS, temel2018cure, temel2019traffic, prabhushankar2022introspective]. Domain adaptation encompasses strategies to reduce this gap when the target data is limited, for instance by aligning feature distributions between source and target domains [ganin2016domain, chen2019temporal]. Another common technique to adapt to target domains is fine-tuning: a network pretrained on a large, generic dataset (the source) is adapted to a more specialized task (the target) by retraining some or all layers, thereby leveraging learned representations while adjusting to new data characteristics [panigrahi2020survey, yosinski2014transferable].
In seismic segmentation, generalization to new domains is particularly difficult due to variability in acquisition parameters, stratigraphy, frequency content, and noise levels across surveys. This is especially true for fault delineation, where subtle and discontinuous features are easily obscured by processing artifacts or geologic heterogeneity [ercoli2023evidencing, dou2022md]. Models trained on one domain (such as synthetic datasets like FaultSeg3D [wu2019faultSeg]) often fail to transfer effectively to real data from different basins. Recent surveys [mousavi_deep-learning_2022, mousavi_applications_2024] provide comprehensive overviews over the way deep learning approaches in exploration seismology struggle with such domain shifts and emphasize fault segmentation as a particularly brittle task.
To mitigate these challenges, researchers have explored a variety of domain adaptation strategies, including feature-space [sun2016deep] and frequency-space [trinidad2023seismic] alignment, adversarial learning [tsai2018learning], and seismic-style transfer [du2022disentangling]. Fault-specific adaptations include synthetic fault injection to enrich training distributions [wang2022structural], contrastive learning schemes designed for curvilinear structures [kokilepersaud2022volumetric], and self-supervised pretraining frameworks that improve transferability across surveys with limited labels [rs16050922, faultsam]. Transfer learning pipelines typically rely on pretrained encoders (either from natural images [he2016deep, simonyan2014very] or large-scale seismic simulations with domain-specific decoders [dou2022md, wu2019faultSeg]) followed by fine-tuning on limited labeled sections. These strategies must navigate a tradeoff between plasticity and stability: aggressive weight updates are prone to inducing catastrophic forgetting [kirkpatrick2017overcoming], while conservative tuning may fail to capture domain-specific features. As a result, generalization in seismic fault delineation requires careful calibration of both model architecture and adaptation strategy to handle the subtle, spatially sparse structures across diverse geological, acquisition, and imaging settings.
II-D Inferential Behavior
In semantic segmentation, inference entails mapping learned feature representations to discrete pixel‐level predictions [long2015fully]. The fidelity of this mapping depends on the architecture’s ability to aggregate context and fuse features: models with narrow receptive fields may accurately localize edges but miss global structure [ronneberger2015u], while those with extensive context capture coherent regions at the expense of sharp boundaries [chen2017deeplab]. During inference, the bias–variance trade‐off emerges as a tension between boundary precision and region consistency, sometimes mitigated by post‐processing, e.g., Conditional Random Fields [krahenbuhl2011efficient] or edge‐aware refinement modules [pohlen2017full].
In seismic fault delineation, these inferential tendencies are amplified by the thin, curvilinear nature of faults and the high noise intrinsic to seismic volumes [ercoli2023evidencing]. U-Net’s symmetric encoder–decoder and skip connections excel at preserving local detail, producing crisp fault traces when the signal‐to‐noise ratio is high [ronneberger2015u], but its reliance on local convolutions and fixed strides can fragment continuous faults under heteroskedastic noise [dou2022md]. DeepLab’s [chen2017deeplab] atrous convolutions and Atrous Spatial Pyramid Pooling (ASPP) module gather multi‐scale context, yielding smoother, globally coherent fault masks; yet the dilation patterns can inadvertently merge adjacent non‐fault discontinuities, introducing false positives along stratigraphic horizons. SegFormer’s transformer‐based encoder captures long‐range dependencies and adapts to complex fault geometries enhancing continuity across sections, though its patch‐based attention can produce coarser boundaries if patch size is not carefully chosen [xie2021segformer].
Furthermore, recent works in both seismic [mousavi_deep-learning_2022, mousavi_applications_2024] and general DL [recht2019imagenet, geirhos2020shortcut] domains emphasize that inferential biases are not merely architectural, but also dataset-driven: synthetic datasets may thus encourage smoother, well-connected predictions, while field datasets can induce models to replicate jagged, discontinuous, or crossover-heavy structures (we analyze these trends in depth using the fault characteristic metrics described in Section II-F). New approaches attempt to counteract these biases by introducing geological priors, like fault continuity preservation [faultsam] or curvature-based losses [kokilepersaud2022volumetric]. These studies highlight that inferential behavior in seismic fault delineation is shaped jointly by architectural design and the structural biases present in training data, underscoring the importance of evaluation frameworks that go beyond pixel overlap to account for geological plausibility.
II-E Performance Evaluation
Evaluation metrics for fault interpretation can be broadly grouped into two categories: region-based and distance-based metrics. Region-based metrics quantify the overlap between predicted and reference fault regions, while distance-based metrics measure the geometric discrepancy between their boundaries.
Region-based metrics. The Dice coefficient $D$ measures the spatial overlap between prediction $P$ and ground truth $G$
$D\;=\;\frac{2\,\lvert P\cap G\rvert}{\lvert P\rvert+\lvert G\rvert}$ (1)
Distance-based metrics. The modified Hausdorff distance $HD$ captures the largest average boundary deviation:
$HD=\max\Biggl(\frac{1}{N_{p}}\sum_{p\in P}\min_{g\in G}d(p,g),\frac{1}{N_{g}}\sum_{g\in G}\min_{p\in P}d(p,g)\Biggr)$ (2)
where $d(\cdot,\cdot)$ is the Euclidean distance between two points, and $N_{p}$} and $N_{g}$} are the number of elements in the sets $P$ and $G$, respectively. Additionally, the Bidirectional Chamfer Distance $BCD$ computes the average shortest distance between boundary points in both directions:
$BCD=\frac{1}{N_{p}}\sum_{p\in G}\min_{g\in G}d(p,g)+\frac{1}{N_{g}}\sum_{g\in G}\min_{p\in P}d(p,g)$ (3)
This categorization provides a basis for later discussion in the Results section, where we compare the behaviors of overlap- and distance-focused metrics.
II-F Fault Characteristics Metrics
To characterize fault networks within each dataset, we measure a set of geometric and topological descriptors. These include Length, Curvature, Sinuosity, Segments, and Stepover Density, which correspond to the statistics reported in Table VI.
Length ($L$) [kim2005relationship] is the total length of all fault traces, computed as
$L=\sum_{i=1}^{N_{\text{faults}}}\ell_{i},$ (4)
where $\ell_{i}$ denotes the length of the $i$-th fault and $N_{\text{faults}}$ is the number of fault traces.
Curvature ($\kappa$) [rutter2018geometry] quantifies the local bending of fault traces. At each arc-length position $s$, curvature is defined as
$\kappa(s)=\left|\frac{d\theta}{ds}\right|,$ (5)
where $\theta$ is the tangent orientation along the fault.
Sinuosity ($S$) [hanson2004style] measures how tortuous a fault is, defined as the ratio of its length to the straight-line distance between its endpoints:
$S=\frac{L_{\text{trace}}}{D_{\text{endpoints}}}.$ (6)
Segments ($N_{\text{seg}}$) [schwartz1989fault] is the total number of discrete fault segments observed in the network.
Stepover Density ($D_{\text{stepover}}$) [manighetti2021fault] measures the relative occurrence of stepovers, normalized by fault length:
$D_{\text{stepover}}=\frac{N_{\text{stepover}}}{L},$ (7)
where $N_{\text{stepover}}$ is the total number of stepovers. These descriptors enable dataset-level comparisons of fault geometry and topology.
TABLE I: Comparison of Netherlands F3 and Thebe Seismic Datasets
Parameter Netherlands F3 (North Sea) [terranubis_f3_demo_2023] Thebe (Exmouth Plateau, Australia) [bailey2013geological]
Location Offshore Netherlands, Block F3 NW Shelf of Australia, Carnarvon Basin
Year of Acquisition 1987 2007
Area Coverage $\sim$387 km2 $\sim$1200 km2
Inline × Crossline 651 × 951 $\sim$3174 × $\sim$1803
Bin Size 25 m × 25 m $\sim$25 m × $\sim$25 m (assumed)
Record Length / Sampling 0–1.848 s TWT, 4 ms $\sim$0–4.5 s TWT, 2–4 ms (typical)
Original Purpose Jurassic/Cretaceous hydrocarbon exploration Triassic gas exploration
Public Wells / Logs 4 wells with logs (sonic/gamma; 2 with density) 2 wells (Thebe-1, Thebe-2) with gas discovery
Data Format 3D post-stack time migrated (SEG-Y, OpenDtect) 3D post-stack time migrated (SEG-Y, fault-labeled)
Structural Setting Shallow deltaic shelf, minor faults, salt dome Rifted margin, rotated fault blocks
Main Stratigraphy Miocene–Pliocene clinoforms; deeper Jurassic/Cretaceous Triassic Mungaroo Formation; Jurassic–Cretaceous seal
Reservoir Target Jurassic sandstones; shallow biogenic gas pockets Triassic fluvial sands (Mungaroo Formation)
Data Availability Fully open (TNO/dGB/OpendTect) Public fault-annotated subset; full survey open-file
Key Features Clinoforms, shallow gas, polygonal faults, salt dome Fault block trap, flat spots, complex fault network
Hydrocarbons Shallow biogenic gas (non-commercial) Confirmed dry gas field ($\sim$2–3 Tcf)
Source / Seal Biogenic gas; intraformational shale; Zechstein salt Mungaroo source/reservoir; Muderong Shale seal
III Fault Delineation Datasets
(a) FaultSeg3D
(b) Cracks
(c) Thebe
Figure 3: Visuals from three datasets: (3(a)) synthesized faults in FaultSeg3D, (3(b)) expert labels in CRACKS, and (3(c)) expert labels in Thebe.
(a) Fault-oriented metrics.
(b) Spectral descriptors.
Figure 4: Charaterization of the three considered datasets using (a) fault-oriented metrics and (b) spectral descriptors.
Among all $74$ datasets used for fault delineation [an2023current], only $4$ field datasets (LANDMASS [alaudah2015curvelet, alaudah2018structure, alfarraj2018multiresolution, long2018comparative, long2015characterization, alaudah2018learning, aribido2020self, aribido2021self], GSB [di2019improving, di2018seismic, an2021deep], Thebe [an2021deep, an2020overlap], CRACKS [prabhushankar2024cracks]) and $4$ synthetic datasets (FaultSeg3D [wu2019faultSeg], Bi’s 3D synthetic [bi2021deep], Wu’s 2D SR [lin2022automatic], Pochet’s 2D synthetic [pochet2018seismic]) open-sourced both seismic data and labels. The low ratio of open-source labeled field data hinders the creation of benchmarks for training and evaluation of models.
Besides the lack of public availability, the different characteristics of the datasets pose challenges to the development of generalizable DL models. Specifically, LANDMASS contains image-level fault labels that cannot be used to numerically evaluate the delineation of pixel-wise faults. GSB contains pixel-wise fault labels annotated on only $5$ crossline sections, which limits the evaluation scalability on large test data. Additionally, it is challenging to achieve generalization using only a small number of labels for finetuning [mustafa2024visual]. In contrast, Thebe provides a large amount of pixel-level geophysicist labels across $1803$ crossline sections. CRACKS provides fault labels of varying quality collected from a group of interpreters, including a geophysicist expert, across $400$ inline sections.
Among the four publicly available datasets, we select Thebe and CRACKS considering the model development and evaluation at the pixel level. All the seismic sections in Pochet’s 2D synthetic contain only one straight fault crossing the entire section, presenting less diversity in angle and density compared to the faults in FaultSeg3D. Both Bi’s 3D synthetic and Wu’s 2D SR [lin2022automatic] originate from FaultSeg3D using the same synthesizing workflow. Thus, we use FaultSeg3D as a reference synthetic dataset with a diverse set of faults. We showcase the acquisition and geological properties of CRACKS and Thebe in Table I.
Given that proprietary seismic data is, by definition, not publicly accessible, our study intentionally focuses on open-source datasets to ensure reproducibility and transparency. This choice aligns with established benchmarking practices in machine learning, where the goal is to provide reproducible and extensible frameworks rather than case-specific proprietary analyses. Despite being open-source, the selected datasets span markedly different geological, statistical, and labeling characteristics, including synthetic versus real data, varying signal-to-noise ratios, and both expert and crowdsourced interpretations. These variations provide meaningful heterogeneity for assessing domain shift and model generalization. To illustrate this diversity, Figure 4(a) presents a comparative visualization of the three datasets using the geometric and topological fault metrics introduced in Section II-F. These metrics quantitatively capture differences in fault density, continuity, and complexity, reinforcing that the chosen datasets represent heterogeneous fault conditions suitable for benchmarking generalization. We also showcase a Uniform Manifold Approximation and Projection (UMAP) embedding of frequency descriptors for all three dataset in Figure 4(b), which reveals three well-separated clusters, confirming systematic acquisition and processing-driven distributional shifts. Below we describe and compare the three datasets used for our benchmarking study in detail.
FaultSeg3D is a 3D synthetic dataset with $220$ volumes each with dimensions of $128\times 128\times 128$ [wu2019faultSeg] . In order to better approximate realistic conditions, the authors added background noise estimated from real seismic volumes. The sampling rate and the frequency of the synthetic data vary across the volume to improve the diversity of the data. An example volume is shown in Fig. 3(a).
CRACKS is an open-source dataset with diverse faults delineated across $400$ inline sections of the Netherlands F3 Block [prabhushankar2024cracks, dgb1987netherlands]. The authors in [alaudah2019machine] open-sourced a fully-annotated 3D geological volume of the Netherlands F3 Block for training different models and comparing the performance with objective metrics. Thus, this volume is one of the most extensively studied geographical zones for developing DL-assisted seismic interpretation frameworks [chowdhury2025unified, benkert2024effective, mustafa2024visual, quesada2024benchmarking, mustafa2024explainable, mustafa2023active, kokilepersaud2022volumetric, benkert2022example, benkert2022reliable, benkert2023samples, zhou2023perceptual, chowdhury2023counterfactual, shafiq2022novel, iqbal2023blind, mustafa2021joint, mustafa2021comparative, mustafa2021man, benkert2021explaining, benkert2021explainable, aribido2021self, aribido2020self, mustafa2020spatiotemporal, mustafa2020joint, soliman2020s, alaudah2019machine, di2019developing, di2019improving, alfarraj2019semisupervised, di2019semi, di2019reflector, alaudah2019facies, mustafa2019estimation, alfarraj2019semi]. The diverse fault features in the F3 block, including major versus minor faults and varying orientations, make it an excellent seismic dataset to train and evaluate fault delineation models [ishak2018application, safari2023structural]. However, the annotations in [alaudah2019machine] do not provide pixel-wise fault labels. Thus, CRACKS open-sources fully hand-labeled fault annotations by a group of $32$ interpreters with varying degrees of expertise and a domain expert geophysicist. This dataset not only establishes a standardized benchmark for objective comparison but also can be used to investigate the impact of multiple sets of labels with varying quality. Three sets of fault annotations are used to investigate the impact of training labels with varying quality, including expert labels and two more sets of lower quality labels from two other annotators. Fig. 3(b) shows an example inline section with expert fault labels from CRACKS. CRACKS and FaultSeg3D share geological similarity in the seismic sections in addition to the similar label density.
Thebe is taken from a seismic survey called Thebe Gas Field in the Exmouth Plateau of the Carnarvan Basin on the NW shelf of Australia [AN2021107219, an2021deep]. The dataset contains $1803$ labeled crossline sections of size $1737\times 3174$, making it the largest publicly-available field dataset. The seismic intensity of this dataset exhibits a low variation/standard deviation, which can be observed from the low contrast in Fig. 3(c).
IV Experimental Setup
The workflow of DL-assisted fault delineation involves multiple choices including ($i$) the decision to fine-tune or not, ($ii$) the selection of the datasets for pre-training and fine-tuning, (iiiiii) the selection of different models, ($iv$) the development of pre-processing and post-processing strategies, and ($v$) the standardization of the evaluation protocols. We summarize these steps in Fig. 5, where the pipeline is organized chronologically in a top-to-bottom fashion to reflect the order in which decisions are made in practice. This systematic layout enables us to holistically compare the effect of each component and their combinations. We provide details on each stage of the pipeline in the remainder of this section.
Figure 5: The block diagram of our experimental setup.
IV-A Data Preparation
IV-A1 Standardizing Fault Annotations using Morphological Operations
There exist significant differences between the thickness of fault annotations in CRACKS, Thebe, and FaultSeg3D. As shown in Fig.6(a), the manually delineated faults in Thebe and CRACKS are considerably thicker than the synthesized faults in FaultSeg3D. Such inconsistency can systematically bias model training and evaluation, especially under pixel-wise loss functions such as the Dice loss. In particular, Dice loss computes similarity between predicted and reference masks by balancing overlap against the number of positive pixels in each mask. Thicker annotations artificially increase the proportion of positive pixels, making the loss less sensitive to small spatial deviations and allowing models to achieve high scores without precisely localizing the fault centerline. Conversely, thinner annotations lead to a stronger penalization of misalignments, requiring sharper localization for similar Dice scores.
Without standardization, these annotation thickness differences could confound cross-dataset comparisons in our benchmark: a model evaluated on thicker annotations may appear more accurate than one evaluated on thinner ones, even if their true localization capability differs. To mitigate this bias, we standardize fault thickness across datasets by processing the manually delineated faults in Thebe and CRACKS to match the thinner style of FaultSeg3D. Specifically, we skeletonize the raw annotations and then apply dilation with a rank-3 structural element to produce uniform thickness and close small gaps, as illustrated in Fig.6(b). This preprocessing, shown in the Preprocess block in Fig.5, is implemented with scikit-image and scipy.
IV-A2 Training and Test Splits
To ensure meaningful evaluations, we adopt a consistent splitting strategy across the considered datasets. Specifically, we maximize diversity in the test set while ensuring no overlap with the training data. For each dataset, we select spatially distinct subsets to prevent redundancies and simulate deployment conditions on previously unseen segments. In CRACKS, which consists of 400 contiguous inlines, we designate the first 30 and last 30 sections as the test set, totaling 60 sections. The remaining 340 central sections are used for training. This split captures geological variability across the volume while maintaining spatial separation between training and test data. For Thebe, we use 400 sections for training and reserve 100 sections from other parts of the volume for testing, following a similar diversity-maximizing approach. For FaultSeg3D, we follow the established setup from [wu2019faultSeg], using 200 synthetic volumes for training and 20 distinct volumes for testing.
(a) Fault labels before morphological operations
(b) Fault labels after morphological operations
Figure 6: Example $128\times 128$ patches of fault annotations with varying thickness in the three datasets. Left: Thebe, middle: CRACKS, right: FaultSeg3D
IV-B Model setup
We consider eight segmentation architectures that are among the most widely used in the seismic interpretation literature [an2021deep, lin2022automatic, rs16050922, 10042020]. This selection was designed to cover the set of architectures commonly adopted for fault delineation, spanning both classical convolutional models and more recent transformer-based variants. Specifically, we include:
  • deeplab: Deeplab [chen2018encoder] architecture with a Resnet50 [he2016deep] backbone
  • deeplab-m:Deeplab architecture with a Mobilenet [sandler2018mobilenetv2] backbone
  • hed: Hollistically-nested edge detection model [xie2015holistically]
  • rcf: Richer convolutional features [liu2017richer]
  • unet: Unet [ronneberger2015u] architecture with a Resnet50 backbone
  • unet++: Unet++ [zhou2018unet++] architecture with a Resnet50 backbone
  • unet-b: Original Unet architecture as presented in [ronneberger2015u]
  • segformer: Transformer-based model introduced in [xie2021segformer]
For each of these models, we perform all pairwise combinations of pretraining-finetuning settings between the $5$ sets of labeled training data in the top left of Fig. 5. For example, a model is first pretrained on CRACKS-expert and then finetuned on all $5$ datasets, resulting in $25$ sets of weights for each considered model. Each of these model weights is then evaluated on CRACKS, Thebe, and FaultSeg3D. To complete our baselines, we also use four ImageNet-pretrained models and finetune them on our seismic datasets.
IV-C Evaluation Setup
IV-C1 Post-processing for Model Predictions
Common practices of evaluating deep fault delineation networks involve two steps [mustafa2024visual, LI2023105412, an2020overlap, an2021deep]: $(i)$ thresholding the network predictions to obtain binary outputs, and $(ii)$ comparing the binary outcomes with the fault test labels using a metric. The test labels are processed with the same morphological operations as the training faults in order to achieve consistent fault thickness at the input and output of a model. Consequently, the thresholding in step $(i)$ needs to be adaptive to the model and the data accordingly, followed by the same morphological operations for numerical evaluation. We compute the optimal threshold for the dataset using the Optimal Dataset Scale (ODS) metric [an2023understanding]. For a model, its optimal threshold is computed using the training set, which is then applied to binarize the predictions on the test data, followed by the same morphological operation for evaluation.
IV-C2 Performance Metrics
As mentioned in Section II-E, pixel-based and distance-based metrics capture different aspects of prediction quality, and each can fail to fully reflect the structural accuracy of the predicted faults, making the evaluation of fault delineation methods an inherently challenging task. Our benchmark study thus adopts a holistic approach to assess prediction quality by considering a combination of multiple metrics alongside subjective inspection. In this study we choose to report one pixel-based metric, Dice Coefficient (defined in Eq. (1)) and two distance-based metrics: BCD and the modified Hausdorff (defined in Eq. 3, and Eq. 2, respectively). These metrics have not only been extensively used in the seismic literature for model performance assessment [wu2019faultSeg, guillon2020ground, 576361], but allow for two different evaluation axes: pixel overlap-based and structure-based.
IV-C3 Fault Characteristic Comparisons
While the individual metrics described in Section II-F quantify dataset characteristics in isolation, comparative metrics evaluate the similarity between predicted and ground-truth faults. We describe below the metrics we use in Table VII to extend the individual descriptors into differences, ratios, and statistical comparisons.
Strike Similarity (StrikeSim) measures the similarity of orientation distributions between predicted ($P(\theta)$) and ground truth ($G(\theta)$) faults using cosine similarity:
$\text{StrikeSim}=\frac{\sum_{\theta}P(\theta)G(\theta)}{\sqrt{\sum_{\theta}P(\theta)^{2}}\,\sqrt{\sum_{\theta}G(\theta)^{2}}}.$ (8)
Curvature Metrics compare curvature distributions across predictions and ground truth. These include:
$\Delta\kappa$ $=\big|\overline{\kappa}_{\text{pred}}-\overline{\kappa}_{\text{gt}}\big|,$ (9)
$\text{RMSE}_{\kappa}$ $=\sqrt{\frac{1}{N}\sum_{i=1}^{N}\left(\kappa_{i}^{\text{pred}}-\kappa_{i}^{\text{gt}}\right)^{2}},$ (10)
$\text{Corr}_{\kappa}$ $=\frac{\text{Cov}\!\left(\kappa^{\text{pred}},\kappa^{\text{gt}}\right)}{\sigma_{\kappa^{\text{pred}}}\sigma_{\kappa^{\text{gt}}}}.$ (11)
Sinuosity Metrics evaluate differences in tortuosity:
$\Delta S$ $=\overline{S}_{\text{pred}}-\overline{S}_{\text{gt}},$ (12)
$R_{S}$ $=\frac{\overline{S}_{\text{pred}}}{\overline{S}_{\text{gt}}}.$ (13)
Length Metrics evaluate differences in fault trace length:
$\Delta L$ $=L_{\text{pred}}-L_{\text{gt}},$ (14)
$R_{L}$ $=\frac{L_{\text{pred}}}{L_{\text{gt}}}.$ (15)
Segment Metrics compare the number of interpreted fault segments:
$\Delta N_{\text{seg}}$ $=N_{\text{seg}}^{\text{pred}}-N_{\text{seg}}^{\text{gt}},$ (16)
$R_{N_{\text{seg}}}$ $=\frac{N_{\text{seg}}^{\text{pred}}}{N_{\text{seg}}^{\text{gt}}}.$ (17)
Stepover Metrics compare the relative frequency of stepovers:
$\Delta D_{\text{stepover}}$ $=D_{\text{stepover}}^{\text{pred}}-D_{\text{stepover}}^{\text{gt}},$ (18)
$R_{D_{\text{stepover}}}$ $=\frac{D_{\text{stepover}}^{\text{pred}}}{D_{\text{stepover}}^{\text{gt}}}.$ (19)
By combining individual metrics with their comparative counterparts, we assess not only whether faults are detected, but also whether their structural and geometric properties are faithfully reproduced.
V Results
The results of our benchmarking experiments are analyzed across three different thematic axes: (1) generalizability and transferability, (2) training dynamics, and (3) metric evaluation.
Figure 7: Visual encoding to represent the different models pretrained and finetuned on different datasets. Border color signifies the pretraining data source and the fill color signifies the fine-tuning dataset. Models are represented using various shapes.
V-A Generalization and Transferability
As mentioned in Section I, models are often used to process data from new surveys that can differ from their original training sources. As such, generalization and transferability are critical for reliable deployment and to reduce the labeling overhead. Understanding how various training regimes perform across domains is key to developing scalable workflows.
In this subsection, the generalization of models trained under different pretraining-finetuning regimes is analyzed across CRACKS, FaultSeg3D, and Thebe. Unless stated otherwise, we structure our scatter plots using the convention depicted in Fig. 7. The shape of a given point encodes the model used, the border color corresponds to the dataset used for pretraining the model, and the fill color corresponds to the datasets used for finetuning the model. All of our reported figures and plots correspond to models being evaluated in a held-out test partition of one of these 3 datasets.
V-A1 Dataset Alignment and Transfer Trends
Fig. 8 provides a macroscopic view of the model behaviors across the different training setups. On CRACKS (Fig. 8(a)), the top-performing configurations are those pretrained on FaultSeg3D data and fine-tuned on CRACKS. The observation indicates strong geophysical commonalities between FaultSeg3D and CRACKS data, and supports the utility of FaultSeg3D data as a viable pretraining source when real annotations are limited. The fill-color distribution in Fig. 8(a) also establishes a hierarchy of effective fine-tuning datasets for CRACKS: CRACKS ¿ FaultSeg3D ¿ Thebe. The poor finetuning performance on Thebe indicates that it is more distributionally distant from CRACKS than FaultSeg3D data.
Furthermore, when tested on FaultSeg3D data (Fig. 8(b)), models pretrained on CRACKS again outperform those trained from scratch, indicating that the aforementioned alignment is reciprocal. However, for Thebe (Fig. 8(c)), top-performing models are those trained from scratch on Thebe itself. Transferring from either FaultSeg3D or CRACKS results in performance degradation, suggesting that Thebe resides in a distinct feature space.
(a) Test on CRACKS dataset
(b) Test on FaultSeg3D dataset
(c) Test on Thebe dataset
Figure 8: Top: All of out models across our different pretraining and finetuning setups. Bottom: Best performing models for each dataset.
V-A2 Domain Shift and Joint Training
Domain shift is a pressing challenge in seismic DL, particularly when deploying models across surveys with different geological properties. When overlooked, this can lead to brittle models that perform well on training data but fail on new volumes, or models that catastrophically forget features from their original data after finetuning. An illustration of this catastrophic forgetting occurs when models pretrained on CRACKS are fine-tuned on Thebe: despite reasonable performance on Thebe, these models experience dramatic performance degradation when re-evaluated on their original domain. In the case of unet++, for instance, the Dice score on CRACKS drops from 0.34 to 0.12, while the BCD increases twofold, indicating a complete erasure of useful representations, and clear case of catastrophic forgetting. A statistical analysis shows that Thebe has a significantly lower standard deviation (0.124) compared to CRACKS (1.149) and FaultSeg3D (1.052). The contrast and intensity variations in Thebe are lower, while CRACKS and FaultSeg3D have diverse intensity distributions, as shown in Fig. 3. An explanation for this catastrophic forgetting phenomenon is that models trained on Thebe learn from a constrained input range and struggle when tested on datasets with richer intensity distributions. These results suggest that other normalization techniques before training could help reduce these distributional mismatches.
To further explore mitigation strategies for catastrophic forgetting, we conducted a focused case study using Elastic Weight Consolidation (EWC) with the deeplabv3 (ResNet-50 backbone) architecture, when transferring from the synthetic FaultSeg3D dataset to the real CRACKS and Thebe volumes. The EWC regularizer constrains the update of parameters that are important for the source task, aiming to preserve previously acquired knowledge during fine-tuning. As shown in Table IV, the expected behavior holds for the FaultSeg3D to CRACKS setting: performance on the target domain decreases slightly relative to standard fine-tuning, but performance on the source domain improves, indicating partial retention of source information. However, this pattern does not hold for the FaultSeg3D to Thebe case, where performance drops on both domains. This outcome is consistent with our broader observations that Thebe represents a strong outlier whose distribution is sufficiently distinct that attempts to retain source-domain information hinder learning of its features altogether. These results highlight that a driving factor in catastrophic forgetting is the distributional structure of the source-target pair.
The empirical pattern we observe (severe performance loss on the source domain after fine-tuning on Thebe, and the mixed effect of EWC across transfer pairs) can be explained by the interaction of three seismic-specific factors. First, faults are sparse, line-like signals: a network’s predictive capacity depends on a small number of localized, edge-sensitive features. Second, amplitude and contrast statistics differ markedly between datasets (e.g., Thebe’s standard deviation = 0.124 vs. CRACKS = 1.149 and FaultSeg3D = 1.052; see Fig. 3); such compression of dynamic range alters internal activation distributions (batchnorm and convolutional responses), rendering source-trained low-level filters less discriminative. Third, regularization schemes like EWC constrain parameters important for the source via a Fisher-based penalty: this helps when the target is moderately different (it preserves source knowledge while allowing limited adaptation, as in FaultSeg3D to CRACKS), but it can prevent learning altogether when the target is an extreme outlier (as in FaultSeg3D to Thebe), producing simultaneous degradation on both domains. Together these effects make seismic catastrophic forgetting both more likely and more abrupt than in many dense, object-centric vision tasks. Practical mitigations that follow from this analysis potentially include input-level alignment (histogram matching or style transfer), adaptive normalization layers, selective (layer-wise) fine-tuning, rehearsal or pseudo-replay, and capacity-aware choices (larger models or explicit adaptation for small models).
As an additional baseline, experiments on joint training on all possible combinations of datasets in a single training round are conducted using the unet model with a ResNet50 backbone. The results for all unet experiments, including these joint training configurations, are shown in Table II. The results further showcase that features learned from CRACKS and FaultSeg3D data align together, but these learned features are not easily transferable to Thebe. On the other hand, Thebe may act as a regularizer when paired with a large dataset like FaultSeg3D, hurting performance on seen datasets (compared to training them individually) but providing modest generalization to unseen ones: performance on the unseen dataset for this setting (i.e. CRACKS) does not drop as badly as that of the unseen dataset for other joint settings (e.g. FaultSeg3D + CRACKS).
As an additional step to assess the robustness and reliability of our results, we also perform our unet experiments on FaultSeg3D using 5 fold cross validation, which we report on Table III. We can observe that the results in these experiments deviate minimally from the ones we originally reported in Table II, showcasing that our experimental framework is robust against stochasticity.
TABLE II: Performance Metrics (DICE, BCD, Hausdorff) for different training schemes on Unet. All values are rounded to three decimals. Schemes are ranked as best, second best or worst based on satisfying at least two of the three metric criteria, to account for cases where metrics disagree. (blue for highest, pink for second, and gray for the worst.)
Training Configuration Test on FaultSeg3D Test on Cracks Test on Thebe
DICE BCD Hausdorff DICE BCD Hausdorff DICE BCD Hausdorff
Individual Training
Thebe 0.111 (+0.0347 , -0.0380) 18.767 (+33.9692 , -6.4945) 16.095 (+33.6256 , -6.3892) 0.020 (+0.0082 , -0.0077) 69.043 (+17.8047 , -13.3343) 49.012 (+19.7797 , -12.1662) 0.182 (+0.0230 , -0.0158) 52.054 (+7.3546 , -8.1523) 31.615 (+8.2126 , -4.7123)
FaultSeg3D 0.4165 (+0.0473, -0.0717) 9.5632 (+14.6540, -5.2962) 7.1908 (+12.4008, -4.1851) 0.1604 (+0.0479, -0.0445) 27.0161 (+8.7621, -6.1338) 20.6379 (+8.1430, -5.3172) 0.0417 (+0.0085, -0.0081) 232.7213 (+29.4637, -28.2944) 189.9385 (+27.9787, -26.0703)
Cracks 0.012(+0.057 , -0.008) 76.904 (+60.57 , -30.058) 57.798 (+60.57 , -30.058) 0.333 (+0.0574 , -0.0819) 11.330 (+10.8538 , -3.1426) 7.341 (+7.1827 , -2.4529) 0.001 (+0.0013 , -0.0007) 1377.945 (+550.5007 , -578.69) 1022.510 (+222.6350 , -321.1643)
Combined Training
FaultSeg3D + Cracks 0.662 (+0.115 , -0.082) 6.143 (+18.342 , -2.673) 5.282 (+15.365 , -1.890) 0.333 (+0.026 , -0.017) 15.436 (+4.649 , -1.583) 11.597 (+5.264 , -2.119) 0.001 (+0.0012 , -0.0007) 659.708 (+37.582 , -49.906) 430.887 (+31.481, -39.851)
FaultSeg3D + Thebe 0.060 (+0.025 , -0.018) 29.834 (+18.619 , -7.824) 20.151 (+14.253 , -5.342) 0.167 (+0.018 , -0.011) 29.717 (+7.782 , -3.947) 26.136 (+6.492 , -2.371) 0.211 (+0.012 , -0.008) 52.952 (+8.459 , -4.671) 43.348 (+7.935 , -3.649)
Cracks + Thebe 0.038 (+0.0475 , -0.0236) 34.183 (+17.2146 , -9.1272) 22.751 (+15.6639 , -7.0688) 0.195 (+0.039 , -0.0389) 17.8 (+6.04 , -3.61) 13.13 (+6.54 , -3.4) 0.112 (+0.021 , -0.0178) 144.67 (+35.17 , -24.26) 98.25 (+21.13 , -18.07)
All 0.337 (+0.2988 , -0.2512) 28.569 (+55.1384 , -20.0588) 23.674 (+35.9425 , -16.5400) 0.1899 (+0.032 , -0.04) 25.62 (+10.37 , -7.3) 17.88 (+10.7 , -7.3) 0.144 (+0.028 , -0.033) 89.69 (+55.91 , -25.90) 75.85 (+58.49 , -27.09)
Fine-Tuning
Cracks $\rightarrow$ FaultSeg3D 0.667 (+0.152 , -0.015) 4.994 (+20.04 , -2.603) 3.821 (+15.34 , -2.923) 0.028 (+0.0270, -0.0328) 73.433 (+19.3709, -15.0939) 45.378 (+16.5661, -12.8940) 0.027 (+0.0026, -0.0026) 165.315 (+14.13, -10.47) 99.996 (+10.03, -5.91)
Cracks $\rightarrow$ Thebe 0.099 (+0.02, -0.034) 21.959 (+34.82, -7.034) 15.771 (+10.93, -4.034) 0.159 (+0.0270, -0.0328) 26.070 (+8.2453, 6.2880) 20.929 (+7.4229, -5.2930) 0.181 (+0.0172, -0.0132) 53.808 (+11.815, -9.61) 42.169 (+13.56, -12.09)
FaultSeg3D $\rightarrow$ Cracks 0.0825 (+0.1060, -0.0587) 136.2934 (+256.0289, -72.6493) 134.0948 (+255.0063, -71.8545) 0.3751 (+0.0539, -0.0650) 8.8064 (+5.0173, -2.6473) 5.5773 (+3.4140, -2.0301) 0.0193 (+0.0055, -0.0045) 213.6949 (+21.8752, -18.4942) 147.2127 (+14.6880, -11.1500)
FaultSeg3D $\rightarrow$ Thebe 0.0745 (+0.1185, -0.0550) 93.4884 (+91.8717, -43.4160) 84.4202 (+84.7160, -38.6755) 0.1349 (+0.0246, -0.0233) 28.7014 (+7.7890, -6.2796) 23.3679 (+8.1318, -6.1587) 0.1582 (+0.0149, -0.0148) 78.6831 (+48.4932, -25.2388) 51.6490 (+47.2406, -19.6911)
Thebe $\rightarrow$ Cracks 0.019 (+0.0082 , -0.0077) 56.145 (+55.3901 , -18.078) 42.899 (+59.5264 , -15.8497) 0.317 (+0.0454 , -0.0514) 10.896 (+3.7560 , -2.1167) 6.656 (+3.2688 , -1.8525) 0.019 (+0.0052 , -0.0050) 161.131 (+14.8345 , -14.0856) 88.344 (+9.4150 , -9.1093)
Thebe $\rightarrow$ FaultSeg3D 0.526 (+0.1585 , -0.2580) 9.319 (+42.6775 , -6.5765) 8.182 (+40.3980 , -5.9642) 0.053 (+0.0116 , -0.0128) 38.141 (+9.1915 , -5.8445) 34.319 (+9.1316 , -5.7956) 0.037 (+0.0029 , -0.0027) 134.522 (+8.3162 , -6.5321) 122.936 (+8.0421 , -6.2061)
TABLE III: U-Net performance on FaultSeg3D dataset using 5-fold cross-validation
Metric Result
Dice $0.4493\pm 0.0285$
Hausdorff $6.1464\pm 0.5806$
BCD $8.6323\pm 0.7638$
TABLE IV: Results on EWC case study
Training configuration Source Target
Dice Hausdorff BCD Dice Hausdorff BCD
FaultSeg3D $\rightarrow$ Cracks 0.025 35.562 48.598 0.1998 7.5489 12.627
FaultSeg3D $\rightarrow$ Thebe 0.012 148.567 179.181 0.061 77.232 112.493
V-A3 Model Capacity and Transferability
It is generally established in the literature that pretraining on a large dataset can boost the performance of a model even in self-supervised settings [rs16050922]. However, in seismic applications, the benefits of pretraining are critically sensitive to the distributional similarity between the source and target domains—a phenomenon that reflects the strong coupling between geologic variability and model transferability. As shown in our experiments, models pretrained on FaultSeg3D or CRACKS and fine-tuned on Thebe underperform compared to models trained from scratch or initialized from ImageNet weights (Fig. 9). In such distributionally mismatched cases, the standard pretaining-finetuning strategy may hinder rather than help performance.
A factor that also plays a role in modulating generalization is model capacity. Fig. 9 shows that large models like segformer benefit from pretraining on CRACKS, while smaller models (e.g., hed, deeplab-m) generalize better when trained from scratch. The observation suggests that both data alignment and model capacity affect transfer effectiveness. Larger architectures, particularly transformer-based models like segformer, possess greater representational capacity and self-attention mechanisms that capture long-range spatial dependencies, an important property for preserving fault continuity and contextual consistency across sections. These characteristics can potentially explain their stronger cross-domain generalization when pretrained features are well aligned. However, the same capacity also makes them more sensitive to distributional mismatches, leading to potential overfitting to domain-specific amplitude statistics or noise patterns when source and target differ substantially. Smaller-capacity models, by contrast, may act as an implicit regularizer, limiting over-specialization and enabling better zero-shot transfer in mismatched scenarios, as can be observed in Fig. 9.
Furthermore, models respond differently to fine-tuning. When pretraining on FaultSeg3D and fine-tuning on Thebe, models degrade in CRACKS performance (Fig. 10, blue circle) shows that Thebe induces domain shifts that are hard to unlearn. Conversely, even though the finetuning dataset is distributionally closer to the target, Thebe-pretrained segformer and deeplab also degrade on CRACKS after finetuning on FaultSeg3D as shown in Fig.  10 (green circles). The asymmetric behavior highlights the difficulty of finding universally robust pretraining strategies.
[…]
TABLE V: Results on domain adaptation case study
Training configuration FDA DANN Finetuning
Dice Hausdorff Dice Hausdorff Dice Hausdorff
FaultSeg3D $\rightarrow$ Cracks 0.139 48.221 0.156 41.058 0.375 5.577
FaultSeg3D $\rightarrow$ Thebe 0.276 20.277 0.319 15.872 0.158 51.649
Figure 9: Individual Models tested on Thebe. Models with blue edges are pretrained on Thebe without finetuning. While other models are pretrained on different datasets. We show that pretraining on another faults dataset is not beneficial compared to using Imagenet weights or training from randomly initialized models.
Figure 10: Individual models tested on CRACKS. (a) Models pre-trained on the FaultSeg3D data, and fine-tuned on different data. (b) Models pre-trained on the Thebe data, and fine-tuned on different data.
V-B Training Dynamics and Inferential Behavior
In this section we analyze the impact of different training dynamics and model architectures on performance and inferential behavior.
V-B1 Window Size and Loss Function
We evaluate the unet model with a ResNet50 backbone using Dice and Binary Cross-Entropy (BCE) losses, as well as across different window sizes: 96, 128, 256 and 512, or up to the size allowed by the original sections in each dataset. The results for these experiments are shown in Fig. 11, where bigger markers correspond to bigger patch sizes.
We can see that two trends emerge in these experiments. First, when using Dice as a loss, models benefit consistently from larger window sizes, with performance generally improving as the spatial context grows. This is likely because larger patches provide the network with a more complete view of fault structures, enabling better continuity modeling across sections. Second, when using BCE loss, this trend does not hold: performance remains flat or slightly declines with larger windows. This discrepancy stems from the class imbalance inherent to fault delineation [sudre2017generalised]. BCE tends to work best when foreground and background classes are more balanced, whereas Dice loss is explicitly designed to handle imbalance. Using smaller patches effectively “zooms in” on the sparse fault regions, increasing the proportion of fault pixels and improving BCE performance.
From a practical perspective, these results suggest that in real-world seismic interpretation, the optimal windowing standard depends on the chosen loss function and the model’s need for contextual information. For losses that handle imbalance well (e.g., Dice), larger tiles are preferable because they capture longer fault segments and contextual cues that allow for improve spatial continuity. For imbalance-sensitive losses (e.g., BCE), smaller tiles may sometimes be beneficial, though at the cost of reduced global context. Given that Dice consistently benefits from larger windows, and that most modern segmentation pipelines for faults employ class-imbalance aware losses, our choice of using Dice with the largest possible window sizes in our benchmark experiments is both empirically supported and aligned with best practices in seismic fault delineation.
Figure 11: Behavior of DICE and BCE losses under different window sizes
V-B2 Model Nuances in Fault Delineation
We also qualitatively observe that each of the evaluated models presents different nuances in the structure of their fault predictions, irrespective of the pretraining or finetuning strategy used, many of which can be observed in Fig. 15. For example, deeplab tends to produce irregular, stair-like faults, while segformer produces thicker, blob-like faults. unet architectures in general tend to produce thinner faults, with unet++ generating more fragmented ones. These architectural signatures are consistent across training setups and highlight the influence model design choices inherently have in shaping the morphology of predicted faults, which is an important consideration when selecting models for downstream tasks or when interpreting evaluation results beyond numerical metrics.
V-C Metric Robustness and Observability
Due to the high correlation among adjacent sections in a seismic volume, deep learning models tend to generate consistent patterns that vary minimally between neighboring sections. Evaluation metrics respond differently to these subtle variations; some metrics heavily penalize these deviations, while others are more tolerant. The effects of structural variations in fault predictions on the evaluation metrics are investigated.
V-C1 Sensitivity to Visual Structure
Distance-based metrics are generally more tolerant to the structure of the predicted fault. This behavior is illustrated in Fig.  12. Where Fig. 12(a) represents the ground truth, while 12(b) and 12(c) show the predictions of two different models. Although the prediction in Fig. 12(b) appears structurally closer to the ground truth, it receives significantly worse BCD and Hausdorff scores compared to the prediction in Fig. 12(c). Since distance-based metrics do not heavily weigh continuity, the inclusion of a few extra pixels around the fault can improve the BCD score even if those pixels lack proper structural alignment. Notably, these patterns are not anomalies, different models often generate such consistent outputs. Consequently, numerical metrics can be misleading and may display discrepancies between one another.
Figure 12: Example of the structure tolerance in distance-based metrics. (a) shows the ground truth fault annotations. (b) and (c) show predictions from two different models. While (b) appears structurally closer to the ground truth, it receives a significantly worse BCD and Hausdorff but a better Dice score.
V-C2 Fault Sparsity
Another issue with distance-based metrics is that they are designed to measure the quality of a single continuous object in the image, whereas faults can consist of multiple sparse objects. This makes distance-based metrics very sensitive to both the number of faults present in the ground truth and their sparsity. We showcase this issue in Fig.  13, where we consider two cases: one with a few faults and another with many faults. Noise is added to both images, and each is compared against its original version. The case with fewer faults contains many outliers compared to the case with many faults, since the added noise often lies far from any existing fault. As discussed in Section II-E, because BCD is bidirectional, it accounts for each added noise pixel by searching for its closest existing fault. Therefore, having fewer faults results in a much worse BCD score. DICE, while also penalizing noise, is more stable across different sparsity levels.
Figure 13: Sensitivity of distance-based metrics to fault density: Sections with many (a) vs. few faults (b) are evaluated under identical noise. Despite equal noise levels, (b) is penalized more by distance-based metrics, while Dice remains stable.
V-C3 Contradictory Scores and Human Judgement
Although pixel-based metrics are more stable, they remain sensitive to slight pixel shifts. In Fig.  14, we show examples corresponding to two different models. The prediction in Fig. 14(c) looks visually closer to the ground truth, but receives a worse Dice score (0.1325 vs. 0.13421) and significantly better BCD and Hausdorff (40.782 vs. 115.436 and 29.0 vs 72.5). Since the faults in both predictions are poorly structured and spatially misaligned, the Dice coefficient (being overlap-based) penalizes them similarly. In contrast, distance-based metrics such as BCD and Hausdorff distances are more sensitive to the spatial coherence of the predicted faults, hence, they are more tolerant to the structure of the faults. These contradictions imply that some metrics may conflict with visual intuition or downstream utility, and point to the need for context-aware metric selection frameworks.
(a) Ground Truth
(b) BCD:115.43, Hausdorff: 72.5, Dice: 0.1342
(c) BCD:40.782, Hausdorff: 29.0, Dice: 0.1325
Figure 14: (a) Ground truth fault annotations. (b) and (c) display predictions from two different models. Although the prediction in (c) appears visually more aligned with the ground truth than (b), it receives a slightly worse Dice score but substantially better BCD and Hausdorff distance
Figure 15: Predictions of the models pretrained on FaultSeg3D tested CRACKS
VI Analysis Using Fault Characteristic Metrics
TABLE VI: Summary statistics of the three datasets: CRACKS (A), Thebe (B), and FS (C). Values are reported as mean $\pm$ standard deviation.
Dataset Length Curvature Sinuosity Segments Stepover Density
CRACKS 6577.75 $\pm$ 1090.82 0.0177 $\pm$ 0.0038 95.35 $\pm$ 54.30 47.22 $\pm$ 7.61 0.978 $\pm$ 0.0046
Thebe 38457.25 $\pm$ 7676.67 0.0261 $\pm$ 0.0058 13.46 $\pm$ 12.32 10.45 $\pm$ 1.85 0.902 $\pm$ 0.0169
FS 16543.79 $\pm$ 3452.73 0.0369 $\pm$ 0.0081 5.70 $\pm$ 6.32 3.18 $\pm$ 1.10 0.623 $\pm$ 0.207
TABLE VII: Evaluation metrics when models are pretrained on Thebe and tested on Thebe, with or without finetuning on other datasets.
Setup Strike Sim. Curvature Curv. RMSE Curv. Corr Sinuosity $\Delta$ Sinuosity Ratio Length $\Delta$ Length Ratio Segment $\Delta$ Segment Ratio Stepover $\Delta$ Stepover Ratio
Thebe Only 0.907 $\pm$ 0.081 -0.001 $\pm$ 0.006 0.128 $\pm$ 0.014 0.019 $\pm$ 0.054 0.246 $\pm$ 0.140 0.829 $\pm$ 0.068 1715.41 $\pm$ 1044 1.649 $\pm$ 0.543 51.47 $\pm$ 10.04 3.292 $\pm$ 0.455 1.779 $\pm$ 6.669 0.993 $\pm$ 0.061
Thebe→CRACKS 0.756 $\pm$ 0.107 0.012 $\pm$ 0.010 0.174 $\pm$ 0.015 -0.010 $\pm$ 0.044 0.080 $\pm$ 0.152 0.950 $\pm$ 0.090 -1593.73 $\pm$ 925.87 0.560 $\pm$ 0.187 67.08 $\pm$ 9.09 4.001 $\pm$ 0.490 4.770 $\pm$ 0.001 0.989 $\pm$ 0.104
Thebe→FS 0.658 $\pm$ 0.053 0.007 $\pm$ 0.010 0.166 $\pm$ 0.017 -0.003 $\pm$ 0.049 -0.055 $\pm$ 0.183 0.968 $\pm$ 0.114 1141.05 $\pm$ 963.78 0.705 $\pm$ 0.241 90.39 $\pm$ 14.91 5.046 $\pm$ 0.788 0.0004 $\pm$ 0.0001 1.004 $\pm$ 0.024
When evaluating transfer setups where all models were pretrained on Thebe and subsequently finetuned on different datasets before being tested back on Thebe, we observe a clear correspondence between the statistical characteristics of the finetuning dataset (Table VI) and the error patterns in the predictions (Table VII). Training and testing solely on Thebe produces the strongest alignment with the ground truth, as indicated by the highest strike similarity (0.907), reflecting that the model effectively captures the long and continuous fault structures characteristic of Thebe. However, this setup also leads to an overestimation of fault length (length ratio of 1.65) and moderate over-segmentation (segment ratio of 3.29), suggesting that the model tends to exaggerate continuity while artificially fragmenting longer structures.
In contrast, when the Thebe-pretrained model is finetuned on CRACKS and evaluated on Thebe, performance degrades in ways consistent with the CRACKS dataset statistics, which emphasize short, jagged, highly segmented faults with frequent crossovers. Strike similarity drops substantially to 0.756, indicating that the orientation of predictions becomes less aligned with Thebe’s long faults. Length is strongly underestimated (length ratio of 0.56), directly mirroring the shorter average fault lengths in CRACKS. Segmentation increases (segment ratio of 4.00), and the stepover delta rises to 4.77, reflecting the fragmentation and crossover-heavy nature of CRACKS. Moreover, spacing becomes strongly negative, showing that the predictions adopt the denser fault placement bias of CRACKS rather than the sparser spacing of Thebe.
A different type of degradation is observed when finetuning on FaultSeg3D, which is dominated by medium-length, smooth faults with very low segmentation and reduced connectivity. Here, strike similarity falls further to 0.658, the lowest among all setups, demonstrating poor alignment with Thebe’s long continuous faults. Length is again underestimated (ratio of 0.70), in line with FaultSeg3D’s shorter structures. Segmentation is highly exaggerated (segment ratio of 5.05), but stepovers almost vanish (stepover delta close to 0), consistent with FaultSeg3D’s low stepover density. The negative spacing indicates that the model no longer respects Thebe’s distribution of fault separations. Thus, while finetuning on FaultSeg3D removes the crossover effects seen with CRACKS, it introduces extreme fragmentation of Thebe’s continuous faults and produces disconnected structures.
In summary, the in-domain Thebe-only model best preserves structural alignment but exaggerates continuity and segmentation, whereas finetuning on CRACKS injects a bias toward jagged, fragmented, and crossover-rich structures, and finetuning on FaultSeg3D enforces smooth, disconnected, and overly fragmented representations. Both transfer setups degrade performance relative to the in-domain baseline, but in ways that directly reflect the statistical properties of the finetuning datasets. This demonstrates that the dataset-specific fault geometry strongly governs the inductive bias of models, even when pretrained on the same source.
Motivated by these observations, we distill them into a self-contained, general “model pick list” to guide out-of-the-box use (Table IX): first, profile the target dataset by labeling its fault–network statistics as High/Medium/Low (H/M/L) relative to Table VI; then select the row whose triggers match that profile and adopt the corresponding pretrain $\rightarrow$ finetune setup; finally, confirm the choice against cross-setup behavior in Table VII and apply the suggested lightweight post-prediction fixes. This table is intentionally model-agnostic with respect to architecture and emphasizes geometry-aware selection aligning to fault length, segmentation, sinuosity, curvature, stepovers, and spacing so that practitioners can choose the most appropriate setup for their basin characteristics without additional tuning.
VII Conclusion
In this work, we present the first large-scale benchmarking study of seismic fault delineation models under domain shift, spanning more than 200 training configurations across three heterogeneous datasets. Our results show that fine-tuning is generally effective when source and target domains are closely aligned, but becomes brittle under stronger shifts, often leading to catastrophic forgetting. Model capacity also modulates transferability: larger architectures such as Segformer tend to adapt more effectively, while smaller models are more sensitive to mismatch. Domain adaptation methods such as FDA and DANN proved beneficial in highly divergent transfers but sometimes degraded performance in more similar settings, highlighting the risk of negative transfer.
Beyond conventional metrics such as Dice or Hausdorff distance, our geometric and topological analysis demonstrated that models also absorb the structural biases of the datasets they are finetuned on: CRACKS-trained models tended to reproduce short, jagged, crossover-rich structures, while FaultSeg3D-trained models favored smoother but disconnected faults. Even when pretrained on the same source, prediction styles were shaped by the statistical properties of the finetuning dataset. These findings underscore that evaluation should not rely only on pixel-level accuracy but also account for structural plausibility and geological realism.
At the same time, important open questions remain. Future research should focus on designing adaptation methods that remain effective across both mild and severe domain shifts, incorporating geological priors into model architectures to mitigate dataset-specific biases, and developing evaluation protocols that balance quantitative rigor with geological interpretability. Addressing these challenges will be essential to building seismic DL pipelines that are not only accurate but also dependable in real-world interpretation workflows.
TABLE VIII: Key observations
Topic Key Findings
Distributional shift among seismic data The intensity standard deviation of the synthetic FaultSeg3D data is similar to that of CRACKS, while both are very different from Thebe’s. This can be attributed to the discrepancy of seismic features across datasets.
Relationship between CRACKS and FaultSeg3D Both CRACKS and FaultSeg3D data benefit from pretraining on the other, outperforming models trained from scratch in either.
Thebe vs. others Thebe does not benefit significantly from pretraining on other data; training from scratch performs best due to distributional shifts.
Model size relationship with pretraining Larger models like segformer and unet (ResNet50) perform well when pretrained on other datasets and finetuned on FaultSeg3D. Smaller models like rcf and hed degrade in performance with pretraining, indicating a lack of transfer capacity.
Fault density FaultSeg3D and CRACKS have dense faults; Thebe faults are sparse, affecting model prediction density.
Joint CRACKS-FaultSeg3D Combining CRACKS and FaultSeg3D data leads to synergistic features and better results.
Joint training with Thebe Adding Thebe acts as a regularizer: performance drops on original domains but improves generalization.
Domain adaptation FDA and DANN improve large-shift transfers (i.e. FaultSeg3D to Thebe) but degrade performance in aligned domains (i.e. FaultSeg3D to CRACKS).
deeplab behavior Produces jagged or stair-like faults.
segformer behavior Tends to generate thick, blob-like faults.
unet/unet++ behavior unet creates thin faults; unet++ tends to produce fragmented ones.
hed/rcf behavior Less adaptable to Thebe due to fault density mismatch; outputs noisy, distorted shapes.
Loss–context relationship Dice loss benefits from larger window sizes (captures fault continuity), while BCE is more effective with smaller patches due to class imbalance.
Metric structural biases Dice penalizes misshaped faults more, while Hausdorff/BCD may still give high scores due to proximity.
Metric sparsity biases Fewer faults lead to harsher penalty in distance metrics; dense faults often score better.
Fault characteristic transfer Models inherit structural/geometric biases of finetuning dataset. For instance, CRACKS fragmented and crossover faults, while FaultSeg3D induces smoother and disconnected faults.
TABLE IX: Out-of-the-box model pick list based on dataset fault–network statistics (Table VI) and cross-setup behavior (Table VII). Label each metric as High/Medium/Low (H/M/L) relative to Table VI, then select the row that matches; confirm with Table VII.
Dataset profile (vs. Table VI) General pick rule (what the model should do) Setup Suggestion Simple rationale (from Table VII) Quick fixes after prediction
Tortuous & highly segmented with frequent stepovers
Triggers: Sinuosity H, Segments H, Stepover density H, Curvature L–M, Length M
[…]
Long, continuous master faults; sparse stepovers (few segments)
Triggers: Length H, Segments L, Stepover density L, Sinuosity L–M, Curvature M
[…]
Balanced / uncertain regime (no extremes) or new basin
Triggers: All metrics within $\pm 1\sigma$ of Thebe (Table VI) or mixed signals
Start with a neutral, orientation-faithful model (high strike similarity) before specializing. Thebe Only Highest strike similarity and stable curvature; safe default when unsure. Enforce minimum segment length; trim dead-ends; lightly straighten detours.
High curvature and high tortuosity (complex relays)
Triggers: Curvature H, Sinuosity H, Stepover density M, Segments M–H
[…]
Stepover-dominated networks with moderate tortuosity
Triggers: Stepover density H, Sinuosity M, Segments M–H, Length M, Curvature L–M
[…]
Abbreviations. H/M/L: high/medium/low relative to Table VI. Length ratio $L_{\text{pred}}/L_{\text{gt}}$; Sinuosity ratio $S_{\text{pred}}/S_{\text{gt}}$; Stepover ratio $D_{\text{stepover}}^{\text{pred}}/D_{\text{stepover}}^{\text{gt}}$; Strike similarity: cosine similarity of strike histograms (see Table VII). Use Table VI to assign H/M/L triggers; confirm with Table VII (ratios near 11, small differences).
Acknowledgment
This work is supported by ML4Seismic Industry Partners at the Georgia Institute of Technology.